### v1

# 0_home_mhaseliu_Documents_Notes_Research_Multiple_Fourier_Mult_Func_Eq.bib

Abstract
Sawin recently gave an axiomatic characterization of multiple Dirichlet series over the function field $\mathbb{F}_{q}(T)$ and proved their existence by exhibiting the coefficients as trace functions of specific perverse sheaves. However, he did not prove that these series actually converge anywhere, instead treating them as formal power series.
In this paper, we prove that these series do converge in a certain region, and moreover that the functions obtained by analytically continuing them satisfy functional equations.
For convergence, it suffices to obtain bounds on the coefficients, for which we use the decomposition theorem for perverse sheaves, in combination with the Kontsevich moduli space of stable maps to construct a suitable compactification.
For the functional equations, the key identity is a multi-variable generalization of the relationship between a Dirichlet character and its Fourier transform; in the multiple Dirichlet series setting, this uses a density trick for simple perverse sheaves and an explicit formula for intermediate extensions from the complement of a normal crossings divisor.
1   Introduction
A multiple Dirichlet series, roughly speaking, is a multi-variable generalization of the well-studied single-variable Dirichlet series. A usual (single-variable) Dirichlet series is a power series in a single complex variable whose coefficients satisfy multiplicativity relations, whereas a multiple Dirichlet series is a power series in several complex variables whose coefficients satisfy certain twisted multiplicativity relations instead.
The traditional perspective is to require the multiple Dirichlet series to additionally satisfy a group of functional equations. Then, to construct a specific multiple Dirichlet series, twisted multiplicativity allows one to reduce to a local construction of coefficients of prime powers, with the corresponding local generating functions satisfying similar functional equations. Over the function field $\mathbb{F}_{q}(T)$, [chinta_mult] proved there is a local-to-global relationship between these local generating functions and the global multiple Dirichet series. [diaconu_pasol_moduli] moreover showed in the specific setting of quadratic Dirichlet L-series that this local-to-global relationship actually uniquely characterizes the multiple Dirichlet series. In particular, they gave an axiomatic characterization of these specific multiple Dirichlet series (one of the axioms being this local-to-global relationship), and in this setting, [whitehead_multiple] in his thesis was able to establish the functional equations using this local-to-global relationship.
Recently, [sawin_general] massively generalized these ideas by giving an axiomatic characterization of general multiple Dirichlet series over $\mathbb{F}_{q}(T)$, and he proved the existence of such multiple Dirichlet series by expressing the coefficients as trace functions of a perverse sheaf. He also showed that these general multiple Dirichlet series generalize many known examples of multiple Dirichlet series that had appeared in the literature.
Sawin, however, left open whether or not these series are genuine analytic functions, as well as which functional equations are satisfied. In this paper, we use geometric methods to answer these questions.
Fix $q$ to be a power of an odd prime. Let $\mathbb{F}_{q}[T]^{+}$ be the set of monic single-variable polynomials over $\mathbb{F}_{q}$, and let $\mathcal{M}_{d}$ be the subset of degree $d$ monic polynomials. Fix a positive integer $n$, and let $\chi\colon\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ be a non-trivial multiplicative character of order $n$. Let $m$ be a positive integer and $M$ be a symmetric $m\times m$ matrix with integer entries (modulo $n$).
Sawin constructs a multiple Dirichlet series
$L\left(u_{1},\ldots,u_{m};M\right)=\sum_{f_{1},\ldots,f_{m}\in\mathbb{F}_{q}[t]^{+}}a\left(f_{1},\ldots,f_{m};M\right)u_{1}^{\deg f_{1}}\cdots u_{m}^{\deg f_{m}},$
where each “$a$-coefficient” $a\left(f_{1},\ldots,f_{m};M\right)$ arises as the trace of Frobenius acting on the stalk (at the tuple $\left(f_{1},\ldots,f_{m}\right)$ viewed as an element of the moduli space of tuples of monic polynomials $\prod_{i=1}^{m}\mathbb{A}_{\mathbb{F}_{q}}^{d_{i}}$) of a certain perverse sheaf defined by the parameters $M$ and the degrees of the polynomials $f_{1},\ldots,f_{m}$.
Our first main result is as follows:
Theorem 1.
Let $n,m,$ and $s$ be positive integers, and $M$ a symmetric $m\times m$ matrix with coefficients in $\mathbb{Z}/n\mathbb{Z}$. Then, $L\left(u_{1},\ldots,u_{m};M\right)$ is an analytic function with a non-empty region of convergence.
The functional equations relate the series $L\left(u_{1},\ldots,u_{m};M\right)$ to another series, which is modified in two ways: 1. some fudge factors are added to the coefficients, and 2. the matrix $M$ is replaced with another matrix $M^{\prime}$.
Let $\psi$ be the non-trivial additive character $e^{2\pi i\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(-)/p}$ on $\mathbb{F}_{q}$, and let $G\left(\chi,\psi\right)=\sum_{x\in\mathbb{F}_{q}^{\times}}\chi(x)\psi(x)$.
Definition 2.
For non-negative integers $d_{s+1},\ldots,d_{m}\geq 0$ and a list of integers $M_{1,s+1},\ldots,M_{1,m}$, let
$\operatorname{fudge}\left(d_{s+1},\ldots,d_{m};M_{1,s+1},\ldots,M_{1,m}\right)\coloneqq\frac{\chi(-1)^{\sum_{s+1\leq i<j\leq m}d_{i}d_{j}M_{1,i}}(-1)^{\sum_{i\geq s+1}\frac{d_{i}(d_{i}-1)(q-1)}{4}}}{\prod_{i\geq s+1}G\left(\chi^{M_{1,i}},\psi\right)^{d_{i}}}.$
Definition 3.
Let $M$ be a symmetric $m\times m$ matrix with coefficients in $\mathbb{Z}/n\mathbb{Z}$. Define $M^{\prime}$ to be another symmetric $m\times m$ matrix such that
  1. (i)
    $M^{\prime}_{i,j}=M_{i,j}+M_{1,i}+M_{1,j}$ for $j>i\geq s+1,$
  2. (ii)
    $M^{\prime}_{i,i}=M_{i,i}+M_{1,i}+n/2$ for $i\geq s+1$,
  3. (iii)
    $M^{\prime}_{1,i}=-M_{1,i}$ for all $i$,
  4. (iv)
    $M_{i,j}^{\prime}=M_{i,j}$ for $j>i>1$ and $i\leq s$, and
  5. (v)
    $M_{i,i}^{\prime}=M_{i,i}$ for $i\leq s$.
Remark 4.
One should think of $a$-coefficients with matrix $M^{\prime}$ (in comparison to $a$-coefficients with matrix $M$) as playing a similar role to the conjugate of a Dirichlet character. In fact, note that $\left(M^{\prime}\right)^{\prime}=M$.
Define $L_{\operatorname{fudge}}\left(u_{1},\ldots,u_{m};M\right)$ to be a slight variant of $L\left(u_{1},\ldots,u_{m};M\right)$:
$L_{\operatorname{fudge}}\left(u_{1},\ldots,u_{m};M\right)$
$=\sum_{f_{1},\ldots,f_{s}\in\mathbb{F}_{q}[t]^{+}}\sum_{d_{s+1},\ldots,d_{m}}b\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{f_{s+1}\in\mathcal{M}_{d_{s+1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M^{\prime}\right)u_{1}^{d_{1}}\cdots u_{m}^{d_{m}},$
where
$b\left(d_{\geq s+1};M_{1,\geq s+1}\right)=\begin{cases}\frac{1}{q^{\sum_{i=s+1}^{m}d_{i}/2}\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}&\text{if }n\text{ divides }\sum_{i=s+1}^{m}d_{i}M_{1,i},\\ \frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{q^{1+\sum_{i=s+1}^{m}d_{i}/2}\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}&\text{else}.\end{cases}$
Remark 5.
In the special case $n=2$ where $\chi$ is a quadratic character and $q\equiv 1\bmod 4$, we have $M^{\prime}=M$ and $L_{\operatorname{fudge}}$ is simply a change of variables of $L$.
Also, note that permuting the variables of $L\left(u_{1},\ldots,u_{m};M\right)$ is equivalent to permuting the entries of $M$ up to adding a sign into the coefficients of the series, so it suffices to consider functional equations in the first variable.
Our second main result is as follows:
Theorem 6.
Let $n$ be a positive even integer, $m$ and $s$ be positive integers, and $M$ a symmetric $m\times m$ matrix with coefficients in $\mathbb{Z}/n\mathbb{Z}$. Without loss of generality, assume $M_{1,1}=\cdots=M_{1,s}=0$ and $M_{1,s+1},\ldots,M_{1,m}$ are not zero, with $s\geq 1$. Let $\zeta_{n}$ be a primitive complex $n$th root of unity. Then, using the notation above, we have the functional equation
$u_{1}\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)$
$=\left(qu_{1}-1\right)L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots u_{s},q^{1/2}u_{1}u_{s+1},\ldots,q^{1/2}u_{1}u_{m};M\right)$
$\qquad-\frac{qu_{1}+u_{1}-2}{n}\sum_{0\leq j\leq n-1}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}q^{1/2}u_{1}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}q^{1/2}u_{1}u_{m};M\right),$
which is an equality of analytic functions on a domain including the region of convergence of both sides.
Remark 7.
The assumption on $n$ is mainly due to the fact that the unique multiplicative character $\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ of order two is simply $\chi^{n/2}$. Note that if $\chi$ has odd order and $q$ is not a power of 2, we can replace $\chi$ with its square root, double $n$, and double the entries of $M$ to get an equivalent series where $n$ is even.
Remark 8.
We are not able to prove that $L\left(u_{1},\ldots,u_{m};M\right)$ has meromorphic continuation to $\mathbb{C}^{m}$ in general, since we only have a subset of the group of possible functional equations (even then, it may not be possible). For this reason, we do not attempt to optimize the region for which $L\left(u_{1},\ldots,u_{m};M\right)$ has meromorphic continuation, i.e. such as using tools like Bochner’s tube theorem (c.f. [BFG]).
We prove the main results in two steps. The first step is to establish the functional equation as a formal equality of power series. The second step is to prove both sides of the functional equation converge in regions of $\mathbb{C}^{m}$, say, $R_{1}$ and $R_{2}$, and to verify that $R_{1}\cap R_{2}$ is not empty. In particular, we prove theorem 1 in the second step and theorem 6 as a combination of both steps.
The first step generalizes proposition 4.4 of [sawin_general], which explicitly computes the $a$-coefficient when $M$ is of the form $\begin{bmatrix}0&-1\\ -1&n/2\end{bmatrix}$. The main trick is to establish a multi-variable variant of the relationship between the conjugate of a Dirichlet character and the Fourier transform of the Dirichlet character. Using this input, obtaining the functional equation is mostly a formal, though tedious, calculation, that closely mirrors that of the single-variable setting.
To establish this connection with the Fourier transform in his example, Sawin uses a density argument, where he first proves the claim easily for tuples of monic polynomials that are square-free and pair-wise relatively prime, and then uses properties of perverse sheaves to extend the result to all tuples.
In our more general setting, the claim, even for tuples of monic polynomials that are square-free and pair-wise relatively prime, is not obvious, and we use another geometric idea: The intermediate extension (which is typically a complex in the derived category) of a tame lisse sheaf with finite monodromy on the complement of a normal crossings divisor has an explicit description as a genuine sheaf. Using this, we obtain an explicit formula for the $a$-coefficients for a slightly larger set of tuples. Then, using a boot-strapping argument, we obtain the relationship between an $a$-coefficient with matrix $M^{\prime}$ (the analogous notion of a conjugate Dirichlet character) and the Fourier transform of an $a$-coefficient with matrix $M$.
The second step is mainly about bounding individual $a$-coefficients for fixed tuples $\left(f_{1},\ldots,f_{m}\right)$ and sums of $a$-coefficients over all tuples $\left(f_{1},\dots,f_{m}\right)$ of fixed degrees. Indeed, with these bounds, obtaining the regions of convergence follows from elementary analysis of power series. The general strategy for bounding comes from another geometric idea, namely by compactifying the space of tuples of monic polynomials that are square-free and pair-wise relatively prime using a quotient of the Kontsevich moduli stack ¯$\mathcal{\overline{\mathcal{M}}}_{0,r}\left(\mathbb{P}^{1},1\right)$—the space of stable maps from a genus zero curve to $\mathbb{P}^{1}$ of degree one—by a Young subgroup of the symmetric group $S_{r}$. The decomposition theorem for perverse sheaves allows us to bound the $a$-coefficients in terms of cohomology of this compactification, which, after translating to the situation over characteristic zero, can be done by bounding the number of cells using a combinatorial argument related to counting rooted planar trees.
Finally, unless otherwise stated, a variety is an irreducible, reduced, separated scheme of finite type over a field.
Acknowledgements. First and foremost, I would like to thank my advisor Will Sawin for suggesting this problem; I’m immensely grateful to Will for his tremendous and invaluable guidance, both high-level and technical, throughout every step of this project. I would also like to thank Adrian Diaconu, Anh Trong Nam Hoang, Donggun Lee, Takyiu Liu, Amal Mattoo, Che Shen, and Fan Zhou for helpful discussions about multiple Dirichlet series, Fox-Neuwirth cells, asymptotics, counting trees, and $S_{n}$-representations, and navigating the literature on moduli spaces of maps.
The author was partially supported by National Science Foundation Grant Number DGE-2036197.
2   Preliminaries
The section comprises a collection of four subsections that recall and prove some technical results that may be of independent interest.
The first subsection gives a quick review of the needed function field number theory and explanation of Sawin’s general construction of multiple Dirichlet series.
The second subsection explains in detail an explicit formula for a specific intermediate extension. A priori, this is an abstract object in the derived category of $\ell$-adic sheaves, but the specific situation we work in allows us to view the intermediate extension as a genuine sheaf. This formula is crucial in both steps of establishing the functional equation. For the first step, the formula is one of the key ingredients in proving the relationship between $a$-coefficients and Fourier transforms. For the second step, the formula, in combination with the decomposition theorem, is the key tool to relate $a$-coefficients to cohomology of lisse sheaves on the Kontsevich moduli space ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$.
The third subsection is about proving some (probably weak) bounds on counting different kinds of trees.
The fourth subsection gives bounds for the cohomology of arbitrary lisse sheaves on the Kontsevich moduli space ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ (quotiented by a Young subgroup). The main input is the bounds on tree counts from the previous subsection.
2.1 Notation and Sawin’s construction
We first establish notation from function field analytic number theory and for the rest of the paper, recalling some from the introduction (for more details, look at [sawin_general]):
  1. (i)
    $q$ is a fixed power of an odd prime $p$.
  2. (ii)
    $\ell$ is a fixed prime not equal to $p$, and we fix an isomorphism $\overline{\mathbb{Q}_{\ell}}\cong\mathbb{C}$.
  3. (iii)
    $\mathbb{F}_{q}[T]$ is the ring of polynomials in one variable over $\mathbb{F}_{q}$.
  4. (iv)
    $\mathbb{F}_{q}[T]^{+}$ is the subset of monic single-variable polynomials over $\mathbb{F}_{q}$.
  5. (v)
    $\mathcal{M}_{t}$ is the subset of monic single-variable polynomials over $\mathbb{F}_{q}$ of degree $t$.
  6. (vi)
    $\mathcal{P}_{<t}$ is the subset of single-variable polynomials over $\mathbb{F}_{q}$ of degree less than $t$.
  7. (vii)
    $m$ and $n$ are fixed positive integers.
  8. (viii)
    $M$ is a symmetric $m\times m$ matrix with entries in $\mathbb{Z}/n\mathbb{Z}$.
  9. (ix)
    $\chi\colon\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ is a non-trivial multiplicative character of order $n$.
  10. (x)
    $\psi\colon\mathbb{F}_{q}\to\mathbb{C}^{\times}$ is the non-trivial additive character $e^{2\pi i\frac{\operatorname{Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(-)}{p}}.$
  11. (xi)
    $G\left(\chi,\psi\right)$ is the Gauss sum $\sum_{x\in\mathbb{F}_{q}^{\times}}\chi(x)\psi(x).$
  12. (xii)
[…]
  13. (xiii)
    The resultant $\operatorname{Res}\left(f,g\right)$ of $f,g\in\mathbb{F}_{q}[T]$ is defined as the product of values of $f$ at the roots of $g$. In particular, $\operatorname{Res}\left(f,g\right)=0$ iff $f$ and $g$ share a common root.
  14. (xiv)
    The residue symbol $\left(\frac{f}{g}\right)_{\chi}$ is defined as $\chi\left(\operatorname{Res}\left(f,g\right)\right)$. An alternative characterization is given by setting $\left(\frac{f}{g}\right)_{\chi}=0$ for $f,g$ sharing a common factor, setting $\left(\frac{f}{g}\right)_{\chi}=\chi\left(f^{\frac{q^{\deg g}-1}{q-1}}\right)$ (with $f^{\frac{q^{\deg g}-1}{q-1}}$ viewed as an element of $\mathbb{F}_{q}$) for $g$ irreducible, and declaring $\left(\frac{-}{-}\right)_{\chi}$ to be separately multiplicative in both inputs.
  15. (xv)
    The residue $\operatorname{Res}(f)$ of a rational function $f$ is defined as the coefficient of $T^{-1}$ when $f$ is written as a Laurent series.
  16. (xvi)
    $e\left(-\right)$ is the composition $\psi\left(\operatorname{Res}\left(-\right)\right).$
Note that affine space $\mathbb{A}_{\mathbb{F}_{q}}^{d}$ can be viewed as a moduli space for monic polynomials of degree $d$. Indeed, for an $\mathbb{F}_{q}$-algebra $R$, $\mathbb{A}_{\mathbb{F}_{q}}^{d}\left(R\right)=R^{d}=\left\{\left(r_{d-1},\ldots,r_{0}\right):r_{i}\in R\right\}$, which we identify with the set of monic single-variable polynomials over $R$ of degree $d$: $\left\{t^{d}+r_{d-1}t^{d-1}+\cdots+r_{0}:r_{i}\in R\right\}$.
Consequently, for non-negative integers $d_{1},\ldots,d_{m}$, we can view $\prod_{i=1}^{m}\mathbb{A}_{\mathbb{F}_{q}}^{d_{i}}$ as a moduli space for tuples of monic polynomials of fixed degrees $d_{1},\ldots,d_{m}$.
Then, for such $d_{1},\ldots,d_{m}$, define the polynomial function
$F_{d_{1},\ldots,d_{m}}=\prod_{i=1}^{m}\operatorname{Res}\left(f_{i}^{\prime},f_{i}\right)^{M_{i,i}}\prod_{1\leq i<j\leq r}\operatorname{Res}\left(f_{i},f_{j}\right)^{M_{i,j}}$
on $\prod_{i=1}^{m}\mathbb{A}^{d_{i}}$.
Let $U$ be the open subset for which $F_{d_{1},\ldots,d_{m}}$ is invertible. Geometrically, $F_{d_{1},\ldots,d_{m}}$ defines a morphism $\prod_{i=1}^{m}\mathbb{A}^{d_{i}}\to\mathbb{A}^{1}$, and $U$ is simply the preimage of $\mathbb{G}_{m}\subset\mathbb{A}^{1}$. In particular, by abuse of notation, $F_{d_{1},\ldots,d_{m}}$ also defines a morphism $U\to\mathbb{G}_{m}$.
On $\mathbb{G}_{m}$, we have a Kummer sheaf associated to $\chi$, which is perhaps best understood as a one-dimensional representation of the etale fundamental group of $\mathbb{G}_{m}$. Namely, there is a natural surjection $\pi_{1}\left(\mathbb{G}_{m}\right)\twoheadrightarrow\mathbb{F}_{q}^{\times}$ arising from Kummer theory, and the composition $\pi_{1}\left(\mathbb{G}_{m}\right)\twoheadrightarrow\mathbb{F}_{q}^{\times}\overset{\overline{\chi}}{\to}\mathbb{C}^{\times}$ gives a continuous one-dimensional representation of $\pi_{1}\left(\mathbb{G}_{m}\right)$. By the correspondence between such representations and lisse rank one etale sheaves, we obtain the Kummer sheaf $\mathcal{L}_{\chi}$ on $\mathbb{G}_{m}$.
Using $F_{d_{1},\ldots,d_{m}}\colon U\to\mathbb{G}_{m}$, we can pull back $\mathcal{L}_{\chi}$ to $U$, which we denote by $\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)$.
Recall that there is an abelian category of “perverse sheaves” inside the derived category of $\ell$-adic sheaves, which is given by the heart of a certain t-structure, c.f. [bbdg]. There are two important examples of perverse sheaves that will appear in this paper:
  1. (i)
    If $X$ is a smooth variety and $L$ is a lisse sheaf on $X$, then $L\left[\dim X\right]$ is perverse.
  2. (ii)
    If $X$ is a variety, $j\colon U\subset X$ is the inclusion of an open subset, and $A$ is a perverse sheaf on $U$, then there is a perverse sheaf $j_{!*}A$ on $X$, defined as the unique extension of $A$ that has no non-trivial sub-objects or quotients supported on $X\backslash U$.
Combining these two examples, $j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[d_{1}+\cdots+d_{m}\right]\right)$ is a perverse sheaf on $\prod_{i=1}^{m}\mathbb{A}^{d_{i}}$, and Sawin defines
$K_{d_{1},\ldots,d_{m}}=j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[d_{1}+\cdots+d_{m}\right]\right)\left[-d_{1}-\cdots-d_{m}\right],$
shifted up so that generically $K_{d_{1},\ldots,d_{m}}$ agrees with $\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)$.
Then, for a tuple $\left(f_{1},\ldots,f_{m}\right)\in\prod_{i=1}^{m}\mathbb{A}^{d_{i}}$, Sawin defines the $a$-coefficient to be
$a\left(f_{1},\ldots,f_{m};M\right)=\operatorname{Tr}\left(\operatorname{Fr}_{q},\left(K_{d_{1},\ldots,d_{m}}\right)_{\left(f_{1},\ldots,f_{m}\right)}\right),$
where $\operatorname{Fr}_{q}$ is the geometric Frobenius.
Before recalling the main theorem of [sawin_general], we review what Sawin calls “compatible systems of sets of ordered pairs of Weil numbers and integers.”
A Weil number is an algebraic number $\alpha$ such that there exists some $i$ such that for any embedding $\overline{\mathbb{Q}_{\ell}}$ into $\mathbb{C}$, the absolute value of the image of $\alpha$ is $q^{i/2}$.
A set of ordered pairs of Weil numbers and integers is simply a set of ordered pairs $\left(\alpha_{j},c_{j}\right)$ indexed by $j$, such that $\alpha_{j}$ is a Weil number, $c_{j}$ is a non-zero integer, and $\alpha_{j}\neq\alpha_{j^{\prime}}$ for $j\neq j^{\prime}$.
A function $\gamma\left(-,-\right)$ from pairs of a prime power $q$ (of a fixed prime $p$) and a multiplicative character $\chi\colon\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ to Weil numbers is said to be a compatible system of Weil numbers if
$\gamma\left(q^{e},\chi_{e}\right)=\gamma\left(q,\chi\right)^{e}$
A function $J\left(-,-\right)$ from pairs of a prime power $q$ (of a fixed prime $p$) and a multiplicative character $\chi\colon\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ to ordered pairs of Weil numbers and integers is said to be a compatible system of sets of ordered pairs of Weil numbers and integers if $J\left(q,\chi\right)=\left\{\left(\alpha_{j},c_{j}\right)\right\}$ means $J\left(q^{e},\chi_{e}\right)=\left\{\left(\alpha_{j}^{e},c_{j}\right)\right\}$.
Finally, the main theorem of [sawin_general] is as follows:
Theorem 9.
[…]
  1. (i)
    (Twisted multiplicativity) If $\prod_{i=1}^{m}f_{i}$ and $\prod_{i=1}^{m}g_{i}$ are relatively prime, then
    $a\left(f_{1}g_{1},\ldots,f_{m}g_{m};M\right)$
    $=a\left(f_{1},\ldots,f_{m};M\right)a\left(g_{1},\ldots,g_{m};M\right)\prod_{1\leq i\leq m}\left(\frac{f_{i}}{g_{i}}\right)_{\chi}^{M_{i,i}}\left(\frac{g_{i}}{f_{i}}\right)_{\chi}^{M_{i,i}}\prod_{1\leq i<j\leq m}\left(\frac{f_{i}}{g_{j}}\right)_{\chi}^{M_{i,j}}\left(\frac{g_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}.$
  2. (ii)
    $a\left(1,\ldots,1;M\right)=a\left(1,\ldots,1,f,1,\ldots,1;M\right)=1$ for all linear polynomials $f$.
  3. (iii)
    $a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)=\left(\frac{\pi^{\prime}}{\pi}\right)_{\chi}^{\sum_{i=1}^{m}e_{i}M_{i,i}}\sum_{j\in J\left(e_{1},\ldots,e_{m};q,\chi,M\right)}c_{j}\alpha_{j}^{\deg\pi}$
    for a prime $\pi$.
  4. (iv)
    $\sum_{f_{1}\in\mathcal{M}_{d_{1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M\right)=\sum_{j\in J\left(d_{1},\ldots,d_{m};q,\chi,M\right)}c_{j}\frac{q^{\sum_{i=1}^{m}d_{i}}}{\overline{\alpha_{j}}}.$
  5. (v)
    $\left|\alpha_{j}\right|\leq q^{\frac{\sum_{i=1}^{m}d_{i}}{2}-1}$ for $\sum_{i=1}^{m}d_{i}\geq 2$.
Existence is given by $a\left(f_{1},\ldots,f_{m};q,\chi,M\right)=a\left(f_{1},\ldots,f_{m};M\right),$ where the right-hand side is the earlier construction in terms of $K_{d_{1},\ldots,d_{m}}$, and $J\left(d_{1},\ldots,d_{m};q,\chi,M\right)$ is the set of ordered pairs of eigenvalues of $\operatorname{Fr}_{q}$ on $\left(K_{d_{1},\ldots,d_{m}}\right)_{\left(t^{d_{1}},\ldots,t^{d_{m}}\right)}$, counted with signed multiplicities (the sign comes from the fact that $\left(K_{d_{1},\ldots,d_{m}}\right)_{\left(t^{d_{1}},\ldots,t^{d_{m}}\right)}$ is a complex).
2.2 An explicit formula for a particular intermediate extension
Proposition 10.
Let $X$ be a smooth variety, $D\hookrightarrow X$ a normal crossings divisor, $U=X\backslash D$, $j\colon U\subset X$ the open immersion, and $L$ a tame lisse sheaf on $U$ with finite monodromy.
Suppose we write $j$ as $j^{\prime}\circ j^{\prime\prime}$, where $j^{\prime\prime}$ is the inclusion of $U$ into the complement of the divisors for which $L$ has trivial local monodromy, and $j^{\prime}$ is the remaining inclusion into $X$.
Then, $j^{\prime\prime}_{*}L$ is a lisse sheaf and
$j_{!*}\left(L[d]\right)[-d]=j^{\prime}_{!}j^{\prime\prime}_{*}L,$
where $d$ is the dimension of $X$.
Proof.
Let us first show that we actually have
$j_{!*}\left(L[d]\right)[-d]=j_{*}L.$
Etale-locally, we may assume $D$ is of the form $t_{1}\cdots t_{s}=0$. Let $X^{\prime}$ be the closed subscheme in $X\times\mathbb{A}^{s}$ cut out by $\left(t^{\prime}_{i}\right)^{N}=t_{i}$ for sufficiently large $N$.
Consider the following Cartesian diagram:
[…]
Here, $\pi$ is the projection onto $X,$ $j$ is the inclusion $U\subset X$, and $\pi^{\prime}$ and $j^{\prime}$ are the respective pull-backs.
By Abhyankar’s lemma (c.f. [Grothendieck_Raynaud]), $\pi^{*}L$ is trivial, $\pi$ is finite, and $\pi^{\prime}$ is etale. Moreover, the monodromy representation of $\pi^{*}L$ factors through an abelian product group $G$, which implies that the representation is the tensor product of one-dimensional representations.
Each of these one-dimensional representations can be realized as monodromy representations associated to Kummer sheaves on $\mathbb{G}_{m}$ because the direct factors of $G$ all look like groups of roots of unity (more precisely inverse limits over $n$ prime to $p$ of $\mu_{n}$).
Since intermediate extension and pushforwards commute with box products, we can assume that $L$ is a lisse rank one sheaf on $\mathbb{G}_{m}$ that has non-trivial local monodromy around the single closed point $\{p\}=\mathbb{A}^{1}\backslash\mathbb{G}_{m}$.
Hence, by definition of the intermediate extension, we have $j_{!*}\left(L[1]\right)=j_{*}L[1]$, from which it follows that
$j_{!*}\left(L[d]\right)[-d]=j_{*}L=j^{\prime}_{*}j^{\prime\prime}_{*}L.$
Next, let us show that the natural map $j^{\prime}_{!}\left(j^{\prime\prime}_{*}L\right)\to j^{\prime}_{*}\left(j^{\prime\prime}_{*}L\right)$ is an isomorphism. It suffices to do this on the level of stalks.
Recall that the local monodromy around an irreducible component is obtained by taking a geometric generic point $\overline{\eta}$, and then pulling $L$ back to $\operatorname{Spec}$ of the fraction field $K_{\overline{\eta}}^{\text{sh}}$ of the strictly Henselian local ring $\mathcal{O}_{X,\overline{\eta}}^{\text{sh}}$, i.e. the strict localization $\widetilde{X}_{\overline{\eta}}$.
By Tag 03Q7 of [stacks], the stalk at $p$ of $j_{*}L$ is given by
$H^{0}\left(\widetilde{\mathbb{A}^{1}}_{\overline{p}}\times_{\mathbb{A}^{1}}\mathbb{G}_{m},L\right)$ $=H^{0}\left(\operatorname{Spec}K_{\overline{p}}^{\text{sh}},L\right)$
$=\left(L_{\overline{p}}\right)^{G_{K_{\overline{p}}^{\text{sh}}}},$
where $G_{K_{\overline{p}}^{\text{sh}}}$ is the absolute Galois group of $K_{\overline{p}}^{\text{sh}}.$ But this is zero for $\overline{p}$ such that the local monodromy is non-trivial.
Finally, we show that $j^{\prime\prime}_{*}L$ is lisse. For the rest of the argument, we may assume $j=j^{\prime\prime}$ by removing all divisors around which the local monodromy is non-trivial.
Suppose the divisors are $D_{1},,\ldots,D_{s}$. Let $g$ be the open immersion $U\subset X\backslash\left\{D_{2}\cup\cdots\cup D_{s}\right\}$. Let us first show that $g_{*}L$ is lisse. Suppose $\overline{\eta}$ is a geometric generic point of $D_{1}$. Since $g$ is an open immersion, pulling $g_{*}L$ back to $\widetilde{X}_{\overline{\eta}}$ is the same as pulling back $L$ to $\operatorname{Spec}K_{\overline{\eta}}^{\text{sh}}=\widetilde{X}_{\overline{\eta}}-\overline{\eta}$, then pushing it forward to $\widetilde{X}_{\overline{\eta}}$. By assumption, this is simply the pushforward of $\overline{\mathbb{Q}_{\ell}}$, so it follows that $g_{*}L$ is lisse on a neighborhood of $\overline{\eta}$. Let $D_{1}^{\prime}$ be the complement of the open locus where $g_{*}L$ is lisse. Since $D_{1}^{\prime}$ does not contain the generic point, it has codimension at least two.
For the sake of contradiction, suppose $D_{1}^{\prime}\neq\emptyset$. Then, let $\overline{\eta^{\prime}}$ be a geometric generic point of $D_{1}^{\prime}$. Again, pulling $g_{*}L$ back to $\widetilde{X}_{\overline{\eta^{\prime}}}$ is the same as pulling $L$ back to $\widetilde{X}_{\overline{\eta^{\prime}}}\times_{X}U=\widetilde{X}_{\overline{\eta^{\prime}}}-\overline{\eta^{\prime}}$, and then pushing forward to $\widetilde{X}_{\overline{\eta^{\prime}}}$. By Grothendieck’s version of Zariski-Nagata purity (c.f. Tag 0BMA of [stacks]), we know $\pi_{1}\left(\widetilde{X}_{\overline{\eta^{\prime}}}-\overline{\eta^{\prime}}\right)=\pi_{1}\left(\widetilde{X}_{\overline{\eta^{\prime}}}\right)=\pi_{1}\left(\overline{\eta^{\prime}}\right)$, which is trivial; indeed, $\widetilde{X}_{\overline{\eta^{\prime}}}-\overline{\eta^{\prime}}$ is an open subset of $\widetilde{X}_{\overline{\eta^{\prime}}}$, so the natural map $\pi_{1}\left(\widetilde{X}_{\overline{\eta^{\prime}}}-\overline{\eta^{\prime}}\right)\to\pi_{1}\left(\widetilde{X}_{\overline{\eta^{\prime}}}\right)$ is surjective, and injectivity follows from the cited statement of purity. Consequently, $g_{*}L$ is lisse in a neighborhood around $\overline{\eta^{\prime}}$, which is a contradiction.
Hence, $g_{*}L$ is lisse. We may then successively apply the same argument to the remaining divisors $D_{2},\ldots,D_{s}$ to conclude that $j_{*}L$ is lisse (because the local monodromy of $g_{*}L$ around each $D_{i}$ is also trivial). ∎
2.3 Elementary combinatorial lemmas on trees
Recall that a rooted tree is simply a tree (an undirected graph that is connected and has no cycles) in which one vertex is designated to be the root. In a rooted tree, a parent of a vertex $v$ is the unique vertex adjacent to $v$ that is on the path to the root; reciprocally, a child of a vertex $v$ is a vertex for which $v$ is its unique parent. The root, in particular, has no parent. A leaf is a vertex with no children. A rooted planar tree is a rooted tree with an embedding in the plane with the root at the top and the children of each vertex $v$ lower than $v$.
Fix $r\geq 1$ and non-negative integers $d_{1},\ldots,d_{m}$ such that $d_{1}+\cdots+d_{m}=r$. Let $S_{r}$ denote the symmetric group on a set of size $r$ and $S_{d_{1}}\times\cdots\times S_{d_{m}}$ be a Young subgroup.
Let $\mathcal{T}$ denote the set of rooted trees with exactly $r$ leaves, up to $S_{d_{1}}\times\cdots\times S_{d_{m}}$-action, such that each non-leaf vertex aside from the root has valence at least three. The leaves are labeled 1 to $r$, and the action is given by $S_{d_{1}}$ acting on the leaves $1$ through $d_{1}$, $S_{d_{2}}$ acting on the leaves labeled $d_{1}+1$ through $d_{2}+d_{1}$, and so on.
For such a tree $T$ and a non-leaf vertex $v$, consider the $d(v)$ (i.e. the valency) neighbors of $v$, including the leaves. If we remove the edge connecting $v$ to a neighbor $w$, then the connected component containing $w$ can be viewed as a rooted tree $T_{w}$ with some number of labels. Each such neighbor defines a new rooted sub-tree with the neighbor as a root. Let $\lambda(v)$ be a partition of $d(v)$ and $S_{\lambda(v)}$ be the product of symmetric groups $S_{\lambda(v)_{1}}\times S_{\lambda(v)_{2}}\times\cdots$ defined as follows:
Consider the trees $T_{w}$ up to isomorphism (for a given $v$), remembering which tree contains the original root (if $v\neq\text{root}$), under the same action. Then, the partition $\lambda(v)$ is a collection of numbers, where each number is the number of $w$ adjacent to $v$ such that $T_{w}$ lies in a particular isomorphism class.
Example 11.
The number of leaves adjacent to $v$ appears in the partition, and if $v\neq\text{root}$, one of the summands of the partition $\lambda(v)$ is simply 1, corresponding to the tree $T_{w}$ which contains the original root.
Lemma 12.
Suppose $T\in\mathcal{T}$.
  1. (i)
    There are at most $r+1$ non-leaf vertices.
  2. (ii)
    The sum of the valences over all non-leaf vertices is at most $3r$.
  3. (iii)
    The sum of the valences over all non-leaf vertices minus the number of non-leaf, non-root vertices is at most $2r$.
  4. (iv)
    The valency of any non-leaf vertex is at most $r+1$.
Proof.
Suppose there are $s$ non-leaf, non-root vertices, whose valencies are $d\left(v_{1}\right),\ldots,d\left(v_{s}\right)$. Then, since the sum of all valencies (including those of the leaves) is twice the number of edges and because a tree has one fewer edge than the total number of vertices, we obtain
[…]
which means
$d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)=r+2s.$
By assumption, the left-hand side is at least $3s$, which means $s\leq r$, i.e. there are at most $r+1$ non-leaf vertices. This also implies $d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)\leq 3r$ and $d\left(v_{1}\right)+\cdots+d\left(v_{s}\right)+d\left(\text{root}\right)-s=r+s\leq 2r$.
Finally, suppose for the sake of contradiction that a non-leaf vertex $v$ has valency at least $r+2$. Also, suppose $v$ has $\ell$ leaves, which is at most $r$. Then, at least $r+2-\ell-1\geq 1$ of these neighbors $w$ are not leaves or the root. By the valency assumption, each $T_{w}$ necessarily contains at least two leaves. Then, $2\left(r+2-\ell-1\right)+\ell$ is at most the total number of leaves, which is $r$. But this implies $\ell\geq r+2$, which is clearly absurd. ∎
Lemma 13.
Let $T\in\mathcal{T}$. There are (${d(\text{root})\choose\lambda(\text{root})}\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash\{1\}}$ distinct rooted planar trees isomorphic to $T$.
Proof.
We prove this inductively. The base case of one vertex is trivial. Starting at the root, say with neighbors $w$, there are (${d(\text{root})\choose\lambda(\text{root})}$ ways to rearrange the different $T_{w}$. Then, recursively, within each $T_{w}$, by induction, there are (${d(w)\choose\lambda(w)}\prod_{v}{d(v)-1\choose\lambda(v)\backslash\{1\}}$ distinct rooted planar trees isomorphic to $T_{w}$, where $v$ runs over all children of $w$ (in particular, $T_{w}$ does not contain the original root). Taking the product over all $w$, the result follows. ∎
Lemma 14.
There are $O_{m}\left(\left(16m\right)^{r}\right)$ distinct rooted planar trees isomorphic to some tree in $\mathcal{T}$.
Proof.
Let $\mathcal{T}^{\prime}$ be the same definition as $\mathcal{T}$, except with $S_{r}$-action replacing $S_{d_{1}}\times\cdots\times S_{d_{m}}$-action. Then, note that the number of trees in $\mathcal{T}$ is at most (${r\choose d_{1},\ldots,d_{m}}$ times the number of trees in $\mathcal{T}^{\prime}$.
We can think of any tree in $\mathcal{T}^{\prime}$ as a rooted unlabelled tree with at least $r+1$ vertices and at most $r+1+r=2r+1$ vertices by lemma 13.
By [Stanley], the number of rooted planar unlabelled trees with $r^{\prime}+1$ vertices is simply the $r^{\prime}$-th Catalan number $C_{r^{\prime}}=\frac{{2r^{\prime}\choose r^{\prime}}}{r^{\prime}+1}$.
Then, the number of distinct rooted planar trees isomorphic to a tree in $\mathcal{T}$ is at most
(${r\choose d_{1},\ldots,d_{m}}\left(C_{r}+\cdots+C_{2r}\right)\leq\frac{r!}{\left(\left(r/m\right)!\right)^{m}}4^{2r}.$
By a variant of Stirling’s approximation, we have
(${r\choose d_{1},\ldots,d_{m}}\left(C_{r}+\cdots+C_{2r}\right)$ $\leq\frac{\sqrt{2\pi r}\left(\frac{r}{e}\right)^{r}e^{\frac{1}{12r}}}{\left(\sqrt{2\pi\frac{r}{m}}\left(\frac{r/m}{e}\right)^{r/m}e^{\frac{1}{12\frac{r}{m}+1}}\right)^{m}}16^{r}$
$=\frac{m^{m/2}e^{\frac{1}{12r}-\frac{m^{2}}{12r+m}}}{\left(\sqrt{2\pi r}\right)^{m-1}}\left(16m\right)^{r}$
$\ll_{m}\left(16m\right)^{r}.$
2.4 Bounds for the cohomology of lisse sheaves on the moduli space of stable maps
Fix a non-negative integer $r.$ Let us quickly recall the definition of the Kontsevich moduli stack ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ and some related objects.
In this subsection, we assume everything is over $\mathbb{C}.$
An $r$-pointed pre-stable curve $\left(C,p_{1},\ldots,p_{r}\right)$ of arithmetic genus zero is a reduced connected projective variety $C$ of dimension one and genus zero with at worst nodal singularities and smooth points $p_{1},\ldots,p_{r}\in C$. A point on a pre-stable curve is special if it is a node or one of the $p_{i}$’s. A stable curve $\left(C,p_{1},\ldots,p_{r}\right)$ of genus zero is a pre-stable curve with at least three special points. A stable map $\left(C,p_{1},\ldots,p_{r}\right)\to\mathbb{P}^{1}$ of degree one is a morphism from an $r$-pointed pre-stable curve such that exactly one of the irreducible components of $C$ maps isomorphically onto $\mathbb{P}^{1}$, the rest of the components are contracted, and each contracted component has at least three special points.
Then, the following are Deligne-Mumford moduli stacks:
  1. (i)
    $\mathcal{M}_{0,r}$ parameterizes $r$-pointed smooth stable curves of genus zero. This space has dimension $r-3$.
  2. (ii)
[…]
  3. (iii)
    $\mathcal{M}_{0,r}\left(\mathbb{P}^{1},1\right)$ parameterizes stable maps $\left(C,p_{1},\ldots,p_{r}\right)\to\mathbb{P}^{1}$ of degree one with $C$ smooth. This space has dimension $r$.
  4. (iv)
    ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ parameterizes stable maps $\left(C,p_{1},\ldots,p_{r}\right)\to\mathbb{P}^{1}$ of degree one. This space has dimension $r$.
[…]
¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}$
denote the preimage of $\mathbb{A}^{r}$ under $\operatorname{ev}$.
Let $\left[-/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]$ denote the stack quotient. We freely use the notation of the previous subsection on trees. For a tree $T\in\mathcal{T}$, let the stabilizer of $T$ under the action of $S_{d_{1}}\times\cdots\times S_{d_{m}}$ be denoted by $\operatorname{stab}(T)$.
Lemma 15.
[¯$\left[\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]$ exhibits a stratification parametrized by trees in $\mathcal{T}$, i.e. of the form
$\mathcal{M}_{T}\coloneqq\left[\left(\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)\times\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{d(\text{root})}}\right)/\operatorname{stab}(T)\right]$
with $T\in\mathcal{T}$ and $v$ runs over all non-leaf and non-root vertices.
Proof.
Recall that the space ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ admits a stratification corresponding to rooted trees with exactly $r$ leaves such that each non-leaf vertex aside from the root has valence at least three, i.e. of the form
$\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)\times\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right).$
Then, ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}$ admits a stratification of the form
$\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)\times\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{d(\text{root})}},$
from which the result follows. ∎
Lemma 16.
Let $f\colon X\to Y$ be a morphism of smooth Deligne-Mumford stacks over a characteristic zero field that is etale-locally on $Y$ isomorphic to a projection map $F\times Y^{\prime}\to Y^{\prime}$. Then, if $\mathcal{L}$ is a lisse sheaf on $X$, the derived pushforward of $\mathcal{L}$, up to shift, is a lisse sheaf.
Proof.
Lisse-ness of $Rf_{*}\mathcal{L}$ is etale-local on $Y$, and since $f$ is etale-locally a projection of the form $F\times Y\to Y$, we may assume $X=F\times Y$. Moreover, in characteristic zero, since $\pi_{1}\left(X\times Y\right)\cong\pi_{1}\left(X\right)\times\pi_{1}\left(Y\right)$, any lisse sheaf on $F\times Y$ can be expressed as the box product of a lisse sheaf on $F$ and a lisse sheaf on $Y$. Then, working etale-locally on $Y$ again, we may assume $\mathcal{L}$ is the pull back of a lisse sheaf from $F$.
By proper base change (c.f. Tag 095S of [stacks]), $Rf_{!}\mathcal{L}$ is isomorphic to the pushforward of a lisse sheaf on $F$ to a point, then pull-backed to $Y$. In other words, $Rf_{!}\mathcal{L}$ is lisse.
If $D$ denotes the Verdier duality functor, then $Rf_{*}\mathcal{L}\cong D\left(Rf_{!}\left(D\left(\mathcal{L}\right)\right)\right)$. Since $X$ is smooth, up to a shift, $D\left(\mathcal{L}\right)$ is lisse. Similarly, since $Y$ is smooth and $Rf_{!}\left(D\left(\mathcal{L}\right)\right)$ is lisse by above, it follows that $D\left(Rf_{!}\left(D\left(\mathcal{L}\right)\right)\right)$ is lisse, up to a shift. ∎
[…]
Proposition 17.
Let $S_{d_{1}}\times\cdots\times S_{d_{m}}$ be a Young subgroup of the symmetric group $S_{r}$. Suppose $U$ is an open substack of ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}$ given by the complement of the union of a (possibly empty) subset of the divisors $D_{i}$. Then, for any local system $\mathcal{L}$ of rank one on $\left[U/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]$, we have
$\dim H^{*}\left(\left[U/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],\mathcal{L}\right)\ll_{m}\left(64m\right)^{r}.$
Proof.
For this proof, when we write $\dim H^{*}(-)$, we mean $\max_{i}\left\{\dim H^{i}\left(-\right)\right\},$ which is well-defined for all situations in consideration.
By lemma 15, [¯$\left[\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]$ admits a stratification corresponding to trees in $\mathcal{T}$, i.e. of the form
$\mathcal{M}_{T}=\left[\left(\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)\times\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{d(\text{root})}}\right)/\operatorname{stab}(T)\right]$
with $T\in\mathcal{T}$.
This means that $\left[U/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]$ admits a stratification of the same form, except with $\mathcal{T}$ replaced with a subset of $\mathcal{T}$.
In particular,
$\dim H^{*}\left(\left[U/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],\mathcal{L}\right)$ $\leq\sum_{T\in\mathcal{T}}\dim H^{*}\left(\mathcal{M}_{T},\mathcal{L}\right).$
For $T\in\mathcal{T}$, note that $\operatorname{stab}(T)$ is a semi-direct product of $S_{\lambda(\text{root})}$ acting on $\left\{\operatorname{stab}(T)\backslash S_{\lambda(\text{root})}\right\}\cup\left\{\operatorname{Id}\right\}$ (the latter being a normal subgroup); indeed, note that the partition $\lambda(v)$ for $v\neq\text{root}$ contains 1 as explained in example 11. Moreover, $\left\{\operatorname{stab}(T)\backslash S_{\lambda(\text{root})}\right\}\cup\left\{\operatorname{Id}\right\}$ is clearly the product of stabilizers of the trees $T_{v}$ for $v$ ranging over non-leaf, non-root vertices. Each such tree is also a rooted tree, so iterating this process implies $\operatorname{stab}(T)$ is an iterated semi-direct product of $S_{\lambda(v)}$’s.
The Leray spectral sequence (c.f. Tag 03QA of [stacks]) applied to the quotient
$q\colon\mathcal{M}_{T}\to\left[\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{d(\text{root})}}/S_{\lambda(\text{root})}\right]$
gives the bound
$\dim H^{*}\left(\mathcal{M}_{T},\mathcal{L}\right)\leq\dim H^{*}\left(\left[\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{d(\text{root})}}/S_{\lambda(\text{root})}\right],Rq_{*}\mathcal{L}\right).$
By lemma 16, $Rq_{*}\mathcal{L}$ is lisse and, by proper base change (c.f. Tag 095S of [stacks]), is of rank at most $\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash S_{\lambda(\text{root})}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],\mathcal{L}\right)$.
Let $\operatorname{PConf}_{d}$ be the configuration space of $d$ points in $\mathbb{C}$. The quotient $\operatorname{PConf}_{d}/S_{\lambda}$ with $\lambda$ a partition of $d$ has a Fox-Neuwirth stratification by Euclidean spaces, whose cells are determined by an ordered partition of $d$ of fixed length, along with an assignment of colors: $\lambda_{1}$ points are colored one color, $\lambda_{2}$ points are colored another, and so on. For more details, c.f. [Fox_Neuwirth].
Since there are (${d\choose\lambda}$ ways to assign colors and at most $2^{d}$ ways to choose an ordered partition of $d$ of fixed length, the number of Fox-Neuwirth cells of $\operatorname{PConf}_{d}\left(\mathbb{C}\right)/S_{\lambda}$ is given by (${d\choose\lambda}2^{d}$.
By definition, $\mathcal{\mathcal{M}}_{0,d(\text{root})}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{d(\text{root})}}$ is the same as $\operatorname{PConf}_{d(\text{root})}$, so it follows that
$\dim H^{*}\left(\mathcal{M}_{T},\mathcal{L}\right)\leq{d(\text{root})\choose\lambda(\text{root})}2^{d(\text{root})}\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash S_{\lambda(\text{root})}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],\mathcal{L}\right).$
Since $\operatorname{stab}(T)$ is an iterated semi-direct product of $S_{\lambda(v)}$’s, we repeat this process using the Leray spectral sequence to obtain
$\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash S_{\lambda(\text{root})}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],\mathcal{L}\right)$
$\leq\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root},w}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash\left\{S_{\lambda(\text{root})}\cup S_{\lambda(w)}\right\}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],R\left(q_{w}\right)_{*}\mathcal{L}\right)$
for some neighbor $w$ of the root (the idea being that $w$ should be thought of as the root of $T_{w}$) and $q_{w}$ the projection map to $\left(\prod_{v\neq\text{root},w}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash\left\{S_{\lambda(\text{root})}\cup S_{\lambda(w)}\right\}\right\}\cup\left\{\operatorname{Id}\right\}\right\}$.
Moreover, by the proof of lemma 7.6 of [EVW], there is an $S_{b}$-equivariant isomorphism
$\operatorname{PConf}_{b}\cong\mathcal{M}_{0,b+1}\times\operatorname{Aff},$
where $\operatorname{Aff}$ is the group scheme of upper triangular matrices corresponding to $z\mapsto Az+B$. The map is given on $S$-points (for an arbitrary scheme $S$) by sending collections of points $p_{1},\ldots,p_{b}\in\mathbb{A}^{1}\left(S\right)$ such that $p_{i}-p_{j}$ is a unit on $S$ for $i\neq j$ to $\left(\left(\mathbb{P}_{S}^{1};\infty,p_{1},p_{2},\ldots,p_{b}\right),\left(p_{2}-p_{1},p_{1}\right)\right)\in\mathcal{M}_{0,b+1}\left(S\right)\times\operatorname{Aff}(S)$. Also, $\operatorname{Aff}\cong\mathbb{G}_{m}\times\mathbb{G}_{a}$.
Let $v$ be a non-leaf and non-root vertex. By example 11, every $\lambda(v)$ contains 1 (as an element of the partition), corresponding to the root. Then, for any rank one local system $\mathcal{L}_{v}$ on $\left[\mathcal{M}_{0,d(v)}/S_{\lambda(v)}\right]$, we have
$\dim H^{*}\left(\left[\mathcal{M}_{0,d(v)}/S_{\lambda(v)}\right],\mathcal{L}_{v}\right)$ $=\dim H^{*}\left(\left[\mathcal{M}_{0,d(v)}/S_{\lambda(v)\backslash 1}\right],\mathcal{L}_{v}\right)$
$\leq\dim H^{*}\left(\left[\operatorname{PConf}_{d(v)-1}/S_{\lambda(v)\backslash 1}\right],\mathcal{L}_{v}\boxtimes\mathbb{Q}_{\ell}\right),$
where in the first line $S_{\lambda(v)\backslash 1}$ acts on the $d(v)-1$ points that do not correspond to the root, and in the second line we use the Kunneth formula (Tag 0F13 of [stacks]).
Then, by applying the same argument about counting Fox-Neuwirth cells (and the lisse-ness of $R\left(q_{w}\right)_{*}\mathcal{L}$), it follows that
$\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root}}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash S_{\lambda(\text{root})}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],\mathcal{L}\right)$
($\leq{d(w)-1\choose\lambda(w)\backslash 1}2^{d(w)-1}\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root},w}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash\left\{S_{\lambda(\text{root})}\cup S_{\lambda(w)}\right\}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],\mathcal{L}\right),$
which means
$\dim H^{*}\left(\mathcal{M}_{T},\mathcal{L}\right)$
($\leq{d(w)-1\choose\lambda(w)\backslash 1}2^{d(w)-1}{d(\text{root})\choose\lambda(\text{root})}2^{d(\text{root})}$
$\qquad\dim H^{*}\left(\left[\left(\prod_{v\neq\text{root},w}\mathcal{M}_{0,d(v)}\right)/\left\{\left\{\operatorname{stab}(T)\backslash\left\{S_{\lambda(\text{root})}\cup S_{\lambda(w)}\right\}\right\}\cup\left\{\operatorname{Id}\right\}\right\}\right],\mathcal{L}\right).$
Iterating this process, it follows that
$\dim H^{*}\left(\mathcal{M}_{T},\mathcal{L}\right)$ $\leq\left(\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash 1}2^{d(v)-1}\right){d(\text{root})\choose\lambda(\text{root})}2^{d(\text{root})}.$
By lemma 12, the sum of valences of non-leaf vertices minus the number of non-leaf, non-root vertices is at most $2r$. As such, we have
$\dim H^{*}\left(\left[U/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],\mathcal{L}\right)$ $\leq\sum_{T\in\mathcal{T}}\dim H^{*}\left(\mathcal{M}_{T},\mathcal{L}\right)$
$\leq\sum_{T\in\mathcal{T}}\left(\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash 1}2^{d(v)-1}\right){d(\text{root})\choose\lambda(\text{root})}2^{d(\text{root})}$
$\leq 2^{2r}\sum_{T\in\mathcal{T}}\left({d(\text{root})\choose\lambda(\text{root})}\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash 1}\right).$
By lemma 13, $\sum_{T\in\mathcal{T}}\left({d(\text{root})\choose\lambda(\text{root})}\prod_{v\neq\text{root}}{d(v)-1\choose\lambda(v)\backslash 1}\right)$ is simply the number of rooted planar trees isomorphic to some tree in $\mathcal{T}$. By lemma 14, this is $O_{m}\left(\left(16m\right)^{r}\right)$.
Then, we have
$\dim H^{*}\left(\left[U/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],\mathcal{L}\right)$ $\ll_{m}4^{r}\cdot\left(16m\right)^{r}$
$=\left(64m\right)^{r},$
as desired.
3   Relationship between $a$-coefficients and Fourier transforms
Let us continue to use the notation from subsection 2.1. Moreover, as in the hypotheses of theorem 6, we assume
  1. (i)
    $n$ is even and
  2. (ii)
    $M_{1,1}=\cdots=M_{1,s}=0$ and $M_{1,s+1},\ldots,M_{1,m}$ are non-zero, with $s\geq 1$.
For a prime polynomial $\pi\in\mathbb{F}_{q}[T]$, a non-trivial multiplicative character $\chi\colon\left(\mathbb{F}_{q}[T]/\pi\right)^{\times}\to\mathbb{C}^{\times}$, a non-trivial additive character $\psi\colon\mathbb{F}_{q}[T]/\pi\to\mathbb{C}^{\times}$, and a polynomial $h\in\mathbb{F}_{q}[T]$ relatively prime to $\pi$, note that
$\sum_{f\in\left(\mathbb{F}_{q}[T]/\pi\right)^{\times}}\chi(f)\psi\left(hf\right)$ $=\sum_{f\in\left(\mathbb{F}_{q}[T]/\pi\right)^{\times}}\chi\left(f/h\right)\psi\left(f\right)$
$=\overline{\chi}\left(h\right)\sum_{f\in\left(\mathbb{F}_{q}[T]/\pi\right)^{\times}}\chi\left(f\right)\psi\left(f\right),$
which gives a relationship between the conjugate Dirichlet character $\overline{\chi}$ and the Fourier transform of $\chi$.
In this section, we establish a variant of this relationship for $a$-coefficients, generalizing a strategy from proposition 4.4 in [sawin_general]. To do so, we simultaneously prove that $a\left(-,f_{2},\ldots,f_{m};M\right)$ is defined independently modulo $f_{s+1}\cdots f_{m}$, which we take to mean that if $f$ and $g$ are monic polynomials differing by a multiple of $f_{s+1}\cdots f_{m}$, then $a\left(f,f_{2},\ldots,f_{m};M\right)=a\left(g,f_{2},\ldots,f_{m};M\right)$.
More precisely, we incrementally prove both claims in special cases that together imply them both in full generality.
Lemma 18.
Whether a divisor $D$ on a variety $X$ is a normal crossings divisors can be checked on closed points.
Proof.
Since the strict Henselization of a local ring preserves (and reflects) regularity, it suffices to show that if for every closed point $p\in D$, $\mathcal{O}_{X,p}$ is regular and there is a regular system of parameters $x_{1},\ldots,x_{d}\in\mathfrak{m}_{p}$ and $1\leq r\leq d$ such that $x_{1}\cdots x_{r}$ cuts out $D$ at $p$, then $D$ is a strict normal crossings divisor (c.f. Tag 0CBN in [stacks]).
Note that in a variety, the set of closed points is dense.
Let $p$ be an arbitrary point of $D$. By Serre’s theorem on openness of regularity, $\mathcal{O}_{X,p}$ is regular. Let $q$ be a closed point that specializes $p$ (i.e. so that $\mathcal{O}_{X,p}$ is a localization of $\mathcal{O}_{X,q}$). Pick a regular system of parameters $x_{1},\ldots,x_{d}\in\mathfrak{m}_{q}$ such that $D$ is cut out by $x_{1}\cdots x_{r}$. Localization is flat, so this regular sequence is mapped to a regular sequence in $\mathcal{O}_{X,p}.$
Lemma 19.
Fix non-negative integers $d_{1},\ldots,d_{m}$ such that $d_{1}\geq d_{s+1}+\cdots+d_{m}$. Let $\left(f_{1},\ldots,f_{m}\right)$ be a closed point of $\prod_{i=1}^{m}\mathbb{A}^{d_{i}}$. Also, let $X\subset\prod_{i=1}^{m}\mathbb{A}^{d_{i}}$ such that $\left(f_{1},f_{1}^{\prime}\right)=\cdots=\left(f_{m},f_{m}^{\prime}\right)=1$ and $\left(f_{i},f_{j}\right)=1$ for $(i,j)\not\in\left\{(1,2),\ldots,(1,m)\right\}$.
Consider the open subset $U\subset X$ such that $\left(f_{1},f_{i}\right)=1$ for all $i\geq s+1$ and $D$ the complement, i.e. $D$ is the locus of points such that $f_{1}$ and some $f_{i}$ share a factor.
Then, $D$ is a normal crossings divisor.
Proof.
We use lemma 18.
Verifying that $D$ is a normal crossings divisor is an etale-local statement, so we may pull back by the factorization/multiplication maps $\underbrace{\mathbb{A}^{1}\times\mathbb{A}^{1}\times\cdots\times\mathbb{A}^{1}}_{d_{s+1}}\to\mathbb{A}^{d_{s+1}},\ldots,\underbrace{\mathbb{A}^{1}\times\mathbb{A}^{1}\times\cdots\times\mathbb{A}^{1}}_{d_{m}}\to\mathbb{A}^{d_{m}},$ which is etale on $X$ by assumption (c.f. the proof of lemma 3.1 in [sawin_general]).
Then, for any $f_{j}(T)$, $D$ etale-locally near $f_{j}(t)$ looks like the union of hyperplanes cut out by $T-\alpha_{ij}$, where the $\alpha_{ij}$’s are the roots of $f_{j}$. For any $S_{j}\subset\{1,2,\ldots,d_{j}\}$, ranging over $s+1\leq j\leq m$, the intersection of the corresponding hyperplanes is
$\left\{f\in\mathbb{A}^{d_{1}}:\prod_{j=s+1}^{m}\prod_{i\in S_{j}}\left(T-\alpha_{ij}\right)|f\right\}\times\mathbb{A}^{d_{2}}\times\cdots\times\mathbb{A}^{d_{s}},$
which has the expected codimension $\left|\bigcup_{j=s+1}^{m}S_{j}\right|$, since $\left|\bigcup_{j=s+1}^{m}S_{j}\right|\leq d_{s+1}+\cdots+d_{m}\leq d_{1}$. ∎
Lemma 20.
Using the notation of lemma 19, let $\left(f_{1},f_{1}^{\prime}\right)=\cdots=\left(f_{m},f_{m}^{\prime}\right)=1$ and $\left(f_{i},f_{j}\right)=1$ for $(i,j)\not\in\left\{(1,2),\ldots,(1,m)\right\}$. Then,
$a\left(f_{1},\ldots,f_{m};M\right)=\prod_{j>i\geq 1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{i\geq s+1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}.$
Proof.
By definition, $a\left(f_{1},\ldots,f_{m};M\right)$ is the trace of Frobenius of the complex
$K_{d_{1},\ldots,d_{m}}=j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[d_{1}+\cdots+d_{m}\right]\right)\left[-d_{1}-\cdots-d_{m}\right].$
Let $j^{\prime}\colon U\subset X$ and $j^{\prime\prime}\colon X\subset\mathbb{A}^{d_{1}}\times\cdots\times\mathbb{A}^{d_{m}}$ as in the notation of lemma 19.
Then, proposition 10 tells us that $K_{d_{1},\ldots,d_{m}}=j^{\prime\prime}_{!*}j^{\prime}_{!}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[d_{1}+\cdots+d_{m}\right]\right)\left[-d_{1}-\cdots-d_{m}\right]$. From this description the result is clear. ∎
Lemma 21.
Suppose $\deg f_{1}\geq\deg f_{s+1}\cdots f_{m}$, i.e. $d_{1}\geq d_{s+1}+\cdots+d_{m}$. Also, assume $\left(f_{i},f_{i}^{\prime}\right)=1$ for $i\geq 1$ and $\left(f_{i},f_{j}\right)=1$ for $(i,j)\not\in\left\{(1,2),\ldots,(1,m)\right\}$. Then,
$a\left(f_{1},\ldots,f_{m};M^{\prime}\right)=\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right).$
Proof.
By lemma 20, we have
$\sum_{h\in\mathcal{M}_{\deg f_{s+1}\cdots f_{m}}}a\left(h,f_{2},\ldots,f_{m},M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right)$
$=\sum_{h\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s}\left(\frac{h}{f_{i}}\right)_{\chi}^{M_{1,i}}\prod_{j>i>1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{i\geq 1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right)$
$=\prod_{j>i>1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{i\geq 1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}\sum_{h\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s+1}\left(\frac{h}{f_{i}}\right)_{\chi}^{M_{1,i}}e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right)$
$=\prod_{j>i>1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{i\geq 1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}\sum_{h\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s+1}\left(\frac{hf_{1}^{-1}}{f_{i}}\right)_{\chi}^{M_{1,i}}e\left(\frac{h}{f_{s+1}\cdots f_{m}}\right)$
$=\prod_{j>i>1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{i\geq 1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}\prod_{i\geq s}\left(\frac{f_{1}}{f_{i}}\right)_{\chi}^{-M_{1,i}}\sum_{h\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s+1}\left(\frac{h}{f_{i}}\right)_{\chi}^{M_{1,i}}e\left(\frac{h}{f_{s+1}\cdots f_{m}}\right).$
Note that $f_{s+1}\cdots f_{m}$ is square-free (since each $f_{i}$ is and no two $f_{i}$’s share a factor). So $f_{s+1}^{\prime}f_{s+2}\cdots f_{m}+\cdots+f_{s+1}\cdots f_{m-1}f_{m}^{\prime}$ is relatively prime to $f_{s+1}\cdots f_{m}$. Note
$\sum_{h\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s+1}\left(\frac{h}{f_{i}}\right)_{\chi}^{M_{1,i}}e\left(\frac{h}{f_{s+1}\cdots f_{m}}\right)$
$=\sum_{h\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s+1}\left(\frac{h}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi\left(\operatorname{Tr}\frac{h}{f_{s+1}^{\prime}f_{s+2}\cdots f_{m}+\cdots+f_{s+1}\cdots f_{m-1}f_{m}^{\prime}}\right),$
since, by the residue theorem, $\operatorname{Res}\left(\frac{h}{f_{s+1}\cdots f_{m}}\right)$ is the sum of residues at the roots $\alpha$ of $f_{s+1}\cdots f_{m}$ (i.e. the coefficients of $\frac{1}{t-\alpha}$ when expressed as a Laurent series), which is the same as the sum over roots $\alpha$ of the $\frac{h}{f_{s+1}^{\prime}f_{s+2}\cdots f_{m}+\cdots+f_{s+1}\cdots f_{m-1}f_{m}^{\prime}}$ evalulated at $\alpha$, which is the trace.
Let $h^{*}=\frac{h}{f_{s+1}^{\prime}f_{s+2}\cdots f_{m}+\cdots+f_{s+1}\cdots f_{m-1}f_{m}^{\prime}}$. Then, the above expression equals
$\sum_{h^{*}\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\prod_{i\geq s+1}\left(\frac{h^{*}\left(f_{s+1}^{\prime}f_{s+2}\cdots f_{m}+\cdots+f_{s+1}\cdots f_{m-1}f_{m}^{\prime}\right)}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi\left(\operatorname{Tr}h^{*}\right)$
$=\sum_{h^{*}\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\left(\frac{h^{*}\left(f_{s+1}^{\prime}f_{s+2}\cdots f_{m}\right)}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{h^{*}\left(f_{s+1}\cdots f_{m-1}f_{m}^{\prime}\right)}{f_{m}}\right)_{\chi}^{M_{1,m}}\psi\left(\operatorname{Tr}h^{*}\right)$
$=\prod_{i\geq s+1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{1,i}}\prod_{i\neq j\geq s+1}\left(\frac{f_{j}}{f_{i}}\right)_{\chi}^{M_{1,i}}\sum_{h^{*}\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\left(\frac{h^{*}}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{h^{*}}{f_{m}}\right)_{\chi}^{M_{1,m}}\psi\left(\operatorname{Tr}h^{*}\right).$
Write $h^{*}=h_{s+1}^{*}f_{s+2}\cdots f_{m}+\cdots+h_{m}^{*}f_{s+1}\cdots f_{m-1}.$ Then
$\sum_{h^{*}\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\left(\frac{h^{*}}{f_{s+1}}\right)^{M_{1,s}}\cdots\left(\frac{h^{*}}{f_{m}}\right)^{M_{1,m}}\psi\left(\operatorname{Tr}h^{*}\right)$
$=\sum_{h^{*}\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}\left(\left(\frac{h_{s+1}^{*}f_{s+2}\cdots f_{m}}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{h_{m}^{*}f_{s+1}\cdots f_{m-1}}{f_{m}}\right)_{\chi}^{M_{1,m}}\right.$
$\qquad\left.\cdot\psi\left(\operatorname{Tr}\left(h_{s+1}^{*}f_{s+2}\cdots f_{m}+\cdots+h_{m}^{*}f_{s+1}\cdots f_{m-1}\right)\right)\right)$
$=\sum_{h_{s+1}^{*}\in\mathbb{F}_{q}[T]/f_{s+1}}\left(\frac{h_{s+1}^{*}}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\psi\left(\operatorname{Tr}h_{s+1}^{*}\right)\cdots\sum_{h_{m}^{*}\in\mathbb{F}_{q}[T]/f_{m}}\left(\frac{h_{m}^{*}}{f_{m}}\right)_{\chi}^{M_{1,m}}\psi\left(\operatorname{Tr}h_{m}^{*}\right)$
$=\prod_{i\geq s+1}G\left(\chi^{M_{1,i}},\psi\right)^{\deg f_{i}}\prod_{i\geq s+1}(-1)^{\frac{\deg f_{i}(\deg f_{i}-1)(q-1)}{4}}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi^{n/2}},$
where in the last line we use the proof of lemma 2.4 of [sawin_general] (and the fact that $n$ is even).
We have
$\sum_{h\in\mathcal{M}_{\deg f_{s+1}\cdots f_{m}}}a\left(h,f_{2},\ldots,f_{m},;M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right)$
$=\prod_{j>i>1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{i=1}^{s}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}\prod_{i\geq s+1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,j}+M_{1,i}+n/2}\prod_{i\geq s+1}\left(\frac{f_{1}}{f_{i}}\right)_{\chi}^{-M_{1,i}}\prod_{i\neq j\geq s+1}\left(\frac{f_{j}}{f_{i}}\right)_{\chi}^{M_{1,i}}$
$\qquad\cdot\prod_{i\geq s+1}G\left(\chi^{M_{1,i}},\psi\right)^{\deg f_{i}}\prod_{i\geq s+1}(-1)^{\frac{\deg f_{i}(\deg f_{i}-1)(q-1)}{4}}$
$=\frac{\prod_{j>i>1,i\leq s}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}}\prod_{j>i\geq s+1}\left(\frac{f_{i}}{f_{j}}\right)_{\chi}^{M_{i,j}+M_{1,i}+M_{1,j}}\prod_{i=1}^{s}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}}\prod_{i\geq s+1}\left(\frac{f_{i}^{\prime}}{f_{i}}\right)_{\chi}^{M_{i,i}+M_{1,i}+n/2}\prod_{i\geq s+1}\left(\frac{f_{1}}{f_{i}}\right)_{\chi}^{-M_{1,i}}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)},$
so the result follows. ∎
Next, we use the density trick in proposition 4.4 of [sawin_general] to remove the assumption “$\left(f_{i},f_{i}^{\prime}\right)=1$ for $i\geq 1$ and $\left(f_{i},f_{j}\right)=1$ for $(i,j)\not\in\left\{(1,2),\ldots,(1,m)\right\}$”. The general idea is to express both sides as trace functions of simple perverse sheaves. Then, by lemma 21, these trace functions agree on a dense open subset, which forces the two perverse sheaves to in fact be the same.
Lemma 22.
Suppose $\deg f_{1}\geq\deg f_{s+1}\cdots f_{m}$, i.e. $d_{1}\geq d_{s+1}+\cdots+d_{m}$. Then,
$a\left(f_{1},\ldots,f_{m};M^{\prime}\right)=\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right).$
Proof.
The argument is more-or-less identical to that of proposition 4.4 of [sawin_general], so we will only sketch the argument. Let $d=d_{s+1}+\cdots+d_{m}$. Recall the $\ell$-adic Fourier transform ([Katz_Laumon]): Let $p_{13}$ and $p_{23}$ be the two projections $\mathbb{A}^{d}\times\mathbb{A}^{d}\times\mathbb{A}^{d}\to\mathbb{A}^{d}\times\mathbb{A}^{d}$ and $\mu\colon\mathbb{A}^{d}\times\mathbb{A}^{d}\times\mathbb{A}^{d}\to\mathbb{A}^{1}$ be the dot product of the first two factors. Then, the Fourier transform $\mathcal{F}_{\psi}\left(-\right)$ is defined as $p_{13!}\left(p_{23}^{*}\left(-\right)\otimes\mu^{*}\mathcal{L}_{\psi}\right)\left[d\right]$.
Let $\sigma\colon\mathbb{A}^{d_{1}}\times\cdots\times\mathbb{A}^{d_{s}}\times\mathbb{A}^{d}\to\mathbb{A}^{d_{1}}\times\cdots\times\mathbb{A}^{d_{s}}\times\mathbb{A}^{d}$ be the morphism sending $\left(f_{1},\ldots,f_{m}\right)$ to
$\left(t^{d}+\operatorname{Res}\left(\frac{T^{d-1}f_{1}}{f_{s+1}\cdots f_{m}}\right)T^{d_{1}-1}+\cdots+\operatorname{Res}\left(\frac{f_{1}}{f_{s+1}\cdots f_{m}}\right),f_{2},\ldots,f_{s},f_{s+1},\ldots,f_{m}\right)$
and $\alpha\colon\mathbb{A}^{d_{1}}\times\cdots\times\mathbb{A}^{d_{s}}\times\mathbb{A}^{d}\to\mathbb{A}^{1}$ is the morphism sending $\left(f_{1},\ldots,f_{m}\right)$ to $\operatorname{Res}\left(\frac{t^{d}f_{1}}{f_{s+1}\cdots f_{m}}\right)$.
Then, the proof of proposition 4.4 of [sawin_general] shows the trace function of
$\sigma^{*}\mathcal{F}_{\psi}K_{d_{1},\ldots,d_{m};M}\otimes\alpha^{*}\mathcal{L}_{\psi},$
where the extra subscript in $K_{d_{1},\ldots,d_{m};M}$ denotes the dependence on the matrix $M$, is
$(-1)^{d}\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right).$
By the Hasse-Davenport relations, after multiplying by a factor of $(-1)^{d}$, $\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)$ is a compatible system of Weil numbers, so there is a lisse rank one sheaf $\mathcal{L}_{G}$ on $\operatorname{Spec}\mathbb{F}_{p}$ such that the trace function of
$\sigma^{*}\mathcal{F}_{\psi}K_{d_{1},\ldots,d_{m};M}\otimes\alpha^{*}\mathcal{L}_{\psi}\otimes\mathcal{L}_{G}$
is
$\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right).$
The proof of proposition 4.4 of [sawin_general] shows that $\sigma^{*}\mathcal{F}_{\psi}K_{d_{1},\ldots,d_{m};M}\otimes\alpha^{*}\mathcal{L}_{\psi}\otimes\mathcal{L}_{G}\left[d_{1}+\cdots+d_{m}\right]$ and $K_{d_{1},\ldots,d_{m};M^{\prime}}\left[d_{1}+\cdots+d_{m}\right]$ are simple perverse sheaves that have trace functions agreeing on the dense open subset defined by $\left(f_{i},f_{i}^{\prime}\right)=1$ for $i\geq 1$ and $\left(f_{i},f_{j}\right)=1$ for $(i,j)\not\in\left\{(1,2),\ldots,(1,m)\right\}$, which, also by the proof of proposition 4.4 of [sawin_general], implies they are isomorphic (in fact, they are both intermediate extensions of the same lisse sheaf). The result follows. ∎
Lemma 23.
Suppose $\deg f_{1},\deg f_{1}^{\prime}\geq\deg f_{s+1}\cdots f_{m}$ such that $f_{1}-f_{1}^{\prime}$ is a multiple of $f_{s+1}\cdots f_{m}$. Then, $a\left(f_{1},f_{2},\ldots,f_{m};M\right)=a\left(f_{1}^{\prime},f_{2},\ldots,f_{m};M\right)$.
Proof.
This follows from lemma 22. Indeed, we have
$a\left(f_{1},\ldots,f_{m};M\right)$ $=\operatorname{fudge}\left(d_{\geq s+1};M^{\prime}_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M^{\prime}\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right)$
$=\operatorname{fudge}\left(d_{\geq s+1};M^{\prime}_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M^{\prime}\right)e\left(\frac{hf_{1}^{\prime}}{f_{s+1}\cdots f_{m}}\right)$
$=a\left(f_{1}^{\prime},\ldots,f_{m};M\right).$
Lemma 24.
Suppose $\left(\frac{-}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{-}{f_{m}}\right)_{\chi}^{M_{1,m}}$ is non-trivial. Then $a\left(-,f_{2},\ldots,f_{m};M\right)$ is independent modulo $f_{s+1}\cdots f_{m}$.
Proof.
Consider the case $a\left(\pi^{a_{1}},\pi^{a_{2}},\ldots,\pi^{a_{m}};M^{\prime}\right)$ and suppose $a_{1}\geq a_{s+1}+\cdots+a_{m}$, where $\pi$ is a prime. Then, by the lemma 22, we have
$a\left(\pi^{a_{1}},\pi^{a_{2}},\ldots,\pi^{a_{m}};M^{\prime}\right)$
$=\operatorname{fudge}\left(a_{\geq s+1}\deg\pi;M_{1,\geq s+1}\right)\cdot\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(h,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right)e\left(\frac{h\pi^{a_{1}}}{\pi^{a_{s+1}+\cdots+a_{m}}}\right)$
$=\operatorname{fudge}\left(a_{\geq s+1}\deg\pi;M_{1,\geq s+1}\right)\cdot\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(h,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right).$
Let $S=\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(h,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right).$ Let $f$ and $\pi$ be relatively prime and such that $\left(\frac{f}{\pi^{a_{s+1}}}\right)_{\chi}^{M_{1,s}}\cdots\left(\frac{f}{\pi^{a_{m}}}\right)_{\chi}^{M_{1,m}}\neq 1$. By lemma 23, note that
$\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(h,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right)=\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(hf,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right)$
because we are summing over all residue classes.
Then, by twisted multiplicativity, we have
$S$ $=\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(h,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right)$
$=\sum_{h\in\mathcal{M}_{\deg\pi^{a_{s+1}+\cdots+a_{m}}}}a\left(hf,\pi^{a_{2}},\ldots,\pi^{a_{m}};M\right)$
$=a\left(f,1,\ldots,1;M\right)\left(\frac{f}{\pi^{a_{s+1}}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{f}{\pi^{a_{m}}}\right)_{\chi}^{M_{1,m}}S$
$=\left(\frac{f}{\pi^{a_{s+1}}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{f}{\pi^{a_{m}}}\right)_{\chi}^{M_{1,m}}S.$
So $S=0$, which means $a\left(\pi^{a_{1}},\pi^{a_{2}},\ldots,\pi^{a_{m}};M^{\prime}\right)=0$.
So, we can assume that for every prime $\pi|f_{1}$, we have $v_{\pi}\left(f_{1}\right)<v_{\pi}\left(f_{s+1}\cdots f_{m}\right),$ where $v_{\pi}\left(f\right)$ denotes the highest power of $\pi$ dividing $f$ (else, $a\left(f_{1},\ldots,f_{m};M\right)=0$ by twisted multiplicativity and the previous case).
Then, by twisted multiplicativity,
$a\left(f_{1}+gf_{s+1}\cdots f_{m},f_{2},\ldots,f_{m};M\right)$
$=a\left(f_{1}\left(1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}\right),f_{2},\ldots,f_{m};M\right)$
$=a\left(f_{1},\ldots,f_{m};M\right)a\left(1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}},1,\ldots,1;M\right)\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{m}}\right)_{\chi}^{M_{1,m}}$
$=a\left(f_{1},\ldots,f_{m};M\right)\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{m}}\right)_{\chi}^{M_{1,m}}.$
Suppose the prime factorization of $f_{s+1}$ is $\pi_{1}^{e_{1}}\cdots\pi_{r}^{e_{r}}$, where each $\pi_{i}$ is a monic prime and $e_{i}$ is a positive integer. Then,
$\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{s+1}}\right)_{\chi}=\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{\pi_{1}}\right)_{\chi}^{e_{1}}\cdots\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{\pi_{r}}\right)_{\chi}^{e_{r}}.$
By assumption, $v_{\pi_{i}}\left(\frac{gf_{s+1}\cdots f_{m}}{f_{1}}\right)\geq 1$ for each $i$, so $\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{s+1}}\right)_{\chi}=1^{e_{1}}\cdots 1^{e_{r}}=1$. Similarly, we have $\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{s+2}}\right)_{\chi}^{M_{1,s+2}},\ldots,\left(\frac{1+\frac{gf_{s+1}\cdots f_{m}}{f_{1}}}{f_{m}}\right)_{\chi}^{M_{1,m}}=1$, so we obtain
$a\left(f_{1}+gf_{s+1}\cdots f_{m},f_{2},\ldots,f_{m};M\right)=a\left(f_{1},\ldots,f_{m};M\right),$
as desired. ∎
Finally, combining the above steps, we are able to prove the relationship in full generality:
Proposition 25.
Let $d_{1},\ldots,d_{m}$ be non-negative integers and $f_{i}\in\mathcal{M}_{d_{i}}$ for each $i$. Then, we have
$a\left(f_{1},\ldots,f_{m};M^{\prime}\right)=\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right).$
Proof.
Pick $v$ relatively prime to $f_{s+1}\cdots f_{m}$ and such that $\deg v+\deg f_{1}\geq\deg f_{s+1}+\cdots+\deg f_{m}.$ Then, by twisted multiplicativity, we have
$a\left(f_{1}v,f_{2},\ldots,f_{m};M^{\prime}\right)$ $=a\left(f_{1},\ldots,f_{m};M^{\prime}\right)a\left(v,1,\ldots,1;M^{\prime}\right)\left(\frac{v}{f_{s+1}}\right)_{\chi}^{-M_{1,s+1}}\cdots\left(\frac{v}{f_{m}}\right)_{\chi}^{-M_{1,m}}$
$=a\left(f_{1},\ldots,f_{m};M^{\prime}\right)\left(\frac{v}{f_{s+1}}\right)_{\chi}^{-M_{1,s+1}}\cdots\left(\frac{v}{f_{m}}\right)_{\chi}^{-M_{1,m}}.$
If $\left(\frac{-}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{-}{f_{m}}\right)_{\chi}^{M_{1,m}}$ is non-trivial, then we can use lemma 24 to see that
$\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}v}{f_{s+1}\cdots f_{m}}\right)=\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(hv^{-1},f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right),$
since we are summing over a full set of residue classes.
Else, by twisted multiplicativity, we have
$a\left(hv^{-1},f_{2},\ldots,f_{m};M\right)$ $=a\left(h,f_{2},\ldots,f_{m};M\right)a\left(v^{-1},1,\ldots,1;M\right)\left(\frac{v^{-1}}{f_{s+1}}\right)_{\chi}^{M_{1,s+1}}\cdots\left(\frac{v^{-1}}{f_{m}}\right)_{\chi}^{M_{1,m}}$
$=a\left(h,f_{2},\ldots,f_{m};M\right),$
so we still have
$\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}v}{f_{s+1}\cdots f_{m}}\right)=\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(hv^{-1},f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right).$
Then, by the special case of lemma 22, we have
$a\left(f_{1}v,f_{2},\ldots,f_{m};M^{\prime}\right)$ $=\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(h,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}v}{f_{s+1}\cdots f_{m}}\right).$
$=\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{h\in\mathcal{M}_{d_{s+1}+\cdots+d_{m}}}a\left(hv^{-1},f_{2},\ldots,f_{m};M\right)e\left(\frac{hf_{1}}{f_{s+1}\cdots f_{m}}\right),$
Since $a\left(hv^{-1},f_{2},\ldots,f_{m};M\right)=a\left(h,f_{2},\ldots,f_{m};M\right)\left(\frac{v}{f_{s+1}}\right)_{\chi}^{-M_{1,s+1}}\cdots\left(\frac{v}{f_{m}}\right)_{\chi}^{-M_{1,m}}$ by the twisted multiplicativity, we get the desired result. ∎
Corollary 26.
$a\left(-,f_{2},\ldots,f_{m};M\right)$ is independent modulo $f_{s+1}\cdots f_{m}$.
4   Derivation of functional equations
We now have all the necessary tools to prove the first step of theorem 6, i.e. the functional equations as a formal equality of power series. Let us continue using the notation from the previous section.
We divide up the analysis into two cases, depending on the triviality of $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}}$ on $\mathbb{F}_{q}^{\times}$ (much like in the single-variable situation). These comprise the subsections 4.1 and 4.2. We then combine these cases to obtain the functional equation in subsection 4.3. Finally, as a quick verification, in subsection 4.4, we show that our functional equations match up with Whitehead’s from his thesis (c.f. [whitehead_multiple]).
For fixed monic polynomials $f_{2},\ldots,f_{m}$ with $\deg f_{i}=d_{i}$, let $d=d_{s+1}+\cdots+d_{m}$. Also, write
$S_{t;f_{2},\ldots,f_{m};M}=\sum_{f\in\mathcal{M}_{t}}a\left(f,f_{2},\ldots,f_{m};M\right)$
and
$P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)=\sum_{t\geq 0}S_{t;f_{2},\ldots,f_{m};M}u_{1}^{t}.$
Lemma 27.
For $0\leq t\leq d$, we have
$S_{t;f_{2},\ldots,f_{m};M}$
$=\frac{q^{t-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[a\left(f_{s}\cdots f_{m},f_{2},\ldots,f_{m};M^{\prime}\right)+\sum_{k=0}^{d-t-2}\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\sum_{f\in\mathcal{M}_{k}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right.$
$\qquad\left.+\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi(-\lambda)\sum_{f\in\mathcal{M}_{d-t-1}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right].$
Proof.
Observe that
$\sum_{f\in\mathcal{M}_{t}}a\left(f,f_{2},\ldots,f_{m};M\right)$ $=\sum_{f\in\mathcal{P}_{<t}}a\left(T^{t}+f,f_{2},\ldots,f_{m};M\right)$
$=q^{t-d}\sum_{f\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}a\left(T^{t}+f,f_{2},\ldots,f_{m};M\right)\sum_{h\in\mathcal{P}_{<d-t}}e\left(\frac{hf}{f_{s+1}\cdots f_{m}}\right)$
$=q^{t-d}\sum_{f\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}a\left(f,f_{2},\ldots,f_{m};M\right)\sum_{h\in\mathcal{P}_{<d-t}}e\left(\frac{hf-hT^{t}}{f_{s+1}\cdots f_{m}}\right)$
$=q^{t-d}\sum_{h\in\mathcal{P}_{<d-t}}e\left(\frac{-hT^{t}}{f_{s+1}\cdots f_{m}}\right)\sum_{f\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}a\left(f,f_{2},\ldots,f_{m};M\right)e\left(\frac{hf}{f_{s+1}\cdots f_{m}}\right).$
Then, using proposition 25, we have
$\frac{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}{q^{t-d}}\sum_{f\in\mathcal{M}_{t}}a\left(f,f_{2},\ldots,f_{m};M\right)$
$=\sum_{h\in\mathcal{P}_{<d-t}}e\left(\frac{-hT^{t}}{f_{s+1}\cdots f_{m}}\right)a\left(h,f_{2},\ldots,f_{m};M^{\prime}\right)$
$=\sum_{h\in\mathcal{P}_{<d-t}}e\left(\frac{-hT^{t}}{f_{s+1}\cdots f_{m}}\right)a\left(h,f_{2},\ldots,f_{m};M^{\prime}\right)$
$=a\left(f_{s+1}\cdots f_{m},f_{2},\ldots,f_{m};M^{\prime}\right)+\sum_{k=0}^{d-t-1}\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\sum_{f\in\mathcal{M}_{k}}e\left(\frac{-\lambda T^{t}f}{f_{s+1}\cdots f_{m}}\right)a\left(\lambda f,f_{2},\ldots,f_{m};M^{\prime}\right)$
$=a\left(f_{s+1}\cdots f_{m},f_{2},\ldots,f_{m};M^{\prime}\right)+\sum_{k=0}^{d-t-1}\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\sum_{f\in\mathcal{M}_{k}}e\left(\frac{-\lambda T^{t}f}{f_{s+1}\cdots f_{m}}\right)a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}$
$=a\left(f_{s+1}\cdots f_{m},f_{2},\ldots,f_{m};M^{\prime}\right)+\sum_{k=0}^{d-t-2}\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1i}}\sum_{f\in\mathcal{M}_{k}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)$
$\qquad+\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi(-\lambda)\sum_{f\in\mathcal{M}_{d-t-1}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right).$
For $t\geq d$, we have $S_{t+1;f_{2},\ldots,f_{m};M}=qS_{t;f_{2},\ldots,f_{m};M}$, so
$S_{d;f_{2},\ldots,f_{m};M}u_{1}^{d}+S_{d+1;f_{2},\ldots,f_{m};M}u_{1}^{d+1}+\cdots$ $=S_{d;f_{2},\ldots,f_{m};M}u^{d}\left(1+qu_{1}+q^{2}u_{1}^{2}+\cdots\right)$
$=\frac{S_{d;f_{2},\ldots,f_{m};M}u_{1}^{d}}{1-qu_{1}}.$
Hence, we have the following lemma.
Lemma 28.
$P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)$ $=\sum_{t=0}^{d-1}S_{t;f_{2},\ldots,f_{m};M}u_{1}^{t}+\frac{S_{d;f_{2},\ldots,f_{m};M}u_{1}^{d}}{1-qu_{1}}.$
Lemma 29.
Let $t\geq 0$. Then,
$qS_{t;f_{2},\ldots,f_{m};M}-S_{t+1;f_{2},\ldots,f_{m};M}$
$=\frac{q^{t+1-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\sum_{f\in\mathcal{M}_{d-t-2}}a(\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right.$
$\qquad\left.+\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi(-\lambda)\left(\sum_{f\in\mathcal{M}_{d-t-1}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)-\sum_{f\in\mathcal{M}_{d-t-2}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right)\right].$
Proof.
For $t\geq d$, both sides are zero by the observation above. For $10\leq t\leq d-$ 1, this follows directly from lemma 27. ∎
We now split our analysis into two cases, depending on whether or not $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}}$ is trivial on $\mathbb{F}_{q}^{\times}$.
4.1 $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}}$ is trivial on $\mathbb{F}_{q}^{\times}$
We have $\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}=1$ and $\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\psi(-\lambda)=-1$.
Lemma 30.
For all $t$, we have
$qS_{t;f_{2},\ldots,f_{m};M}-S_{t+1;f_{2},\ldots,f_{m};M}=\frac{q^{t+1-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[qS_{d-t-2;f_{2},\ldots,f_{m};M^{\prime}}-S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}\right].$
Proof.
If $d=0$, then both sides are 0, so we may assume $d\geq 1$.
For $t\geq 0$, by lemma 29, we have
$qS_{t;f_{2},\ldots,f_{m};M}-S_{t+1;f_{2},\ldots,f_{m};M}$
$=\frac{q^{t+1-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[q\sum_{f\in\mathcal{M}_{d-t-2}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)-\sum_{f\in\mathcal{M}_{d-t-1}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right]$
$=\frac{q^{t+1-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[qS_{d-t-2;f_{2},\ldots,f_{m};M^{\prime}}-S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}\right].$
For $t\leq-1$, since $d-t-2\geq 0$ and $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}^{\prime}}$ is also trivial on $\mathbb{F}_{q}^{\times}$, we obtain
$qS_{d-t-2;f_{2},\ldots,f_{m};M^{\prime}}-S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}=\frac{q^{-t-1}}{\operatorname{fudge}\left(d_{\geq s+1};M^{\prime}_{1,\geq s+1}\right)}\left[qS_{t;f_{2},\ldots,f_{m};M}-S_{t+1;f_{2},\ldots,f_{m};M}\right],$
which implies
$qS_{t;f_{2},\ldots,f_{m};M}-S_{t+1;f_{2},\ldots,f_{m};M}$
$=\frac{q^{t+1-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[qS_{d-t-2;f_{2},\ldots,f_{m};M^{\prime}}-S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}\right],$
as desired. ∎
We then obtain the following functional equation.
Proposition 31.
$P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)\left(qu_{1}-1\right)$ $=\frac{1}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}\left(1-u_{1}\right)P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right).$
Proof.
Using lemmas 28 and 30, we have
$P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)\left(qu_{1}-1\right)$
$=\left(qu_{1}-1\right)\sum_{t=0}^{d-1}S_{t;f_{2},\ldots,f_{m};M}u_{1}^{t}-S_{d;f_{2},\ldots,f_{m};M}u_{1}^{d}$
$=\sum_{t=0}^{d-1}\left(qS_{t;f_{2},\ldots,f_{m};M}u_{1}^{t+1}-S_{t;f_{2},\ldots,f_{m};M}u_{1}^{t}\right)-S_{d;f_{2},\ldots,f_{m};M}u_{1}^{d}$
$=\left(qS_{-1;f_{2},\ldots,f_{m};M}-S_{0;f_{2},\ldots,f_{m};M}\right)+\cdots+\left(qS_{d-1;f_{2},\ldots,f_{m};M}u_{1}^{d}-S_{d;f_{2},\ldots,f_{m};M}u_{1}^{d}\right)$
$=\frac{\left(q^{1-d}S_{d-1;f_{2},\ldots,f_{m};M^{\prime}}-q^{-d}S_{d;f_{2},\ldots,f_{m};M^{\prime}}\right)+\cdots+\left(qS_{-1;f_{2},\ldots,f_{m};M^{\prime}}-S_{0;f_{2},\ldots,f_{m};M^{\prime}}\right)u_{1}^{d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}$
$=\frac{1}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[\left(qS_{-1;f_{2},\ldots,f_{m};M^{\prime}}u_{1}^{d}+\cdots+q^{2-d}S_{d-2;f_{2},\ldots,f_{m};M^{\prime}}u_{1}+q^{1-d}S_{d-1;f_{2},\ldots,f_{m};M^{\prime}}\right)\right.$
$\qquad\left.-\left(S_{0;f_{2},\ldots,f_{m};M^{\prime}}u_{1}^{d}+\cdots+q^{1-d}S_{d-1;f_{2},\ldots,f_{m};M^{\prime}}u_{1}+q^{-d}S_{d;f_{2},\ldots,f_{m};M^{\prime}}\right)\right].$
Also, note that
$u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)$
$=u_{1}^{d-1}\left(\sum_{t=0}^{d-1}S_{t;f_{2},\ldots,f_{m};M^{\prime}}q^{-t}u_{1}^{-t}+\frac{S_{d;f_{2},\ldots,f_{m};M^{\prime}}q^{-d}u_{1}^{-d}}{1-u_{1}^{-1}}\right)$
$=S_{0;f_{2},\ldots,f_{m};M^{\prime}}u_{1}^{d-1}+\cdots+S_{d-2;f_{2},\ldots,f_{m};M^{\prime}}q^{2-d}u_{1}+S_{d-1;f_{2},\ldots,f_{m};M^{\prime}}q^{1-d}+\frac{q^{-d}S_{d;f_{2},\ldots,f_{m};M^{\prime}}}{u_{1}-1},$
so
$P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)\left(qu_{1}-1\right)$ $=\frac{1}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[\left(u_{1}^{d-1}-u_{1}^{d}\right)P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)\right]$
$=\frac{1}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}\left(1-u_{1}\right)P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right),$
as desired. ∎
4.2 $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}}$ is not trivial on $\mathbb{F}_{q}^{\times}$
Note that $\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}=\prod_{i=s+1}^{m}\chi(\lambda)^{d_{i}M_{1,i}}=\chi(\lambda)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}$.
Lemma 32.
For all $t$, we have
$S_{t;f_{2},\ldots,f_{m};M}=\frac{q^{t-d}\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}.$
Proof.
By proposition 25, we have
$S_{d;f_{2},\ldots,f_{m};M}$ $=\sum_{f\in\mathbb{F}_{q}[T]/f_{s+1}\cdots f_{m}}a\left(f,f_{2},\ldots,f_{m};M\right)$
$=\frac{1}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}a\left(f_{s+1}\cdots f_{m},f_{2},\ldots,f_{m};M^{\prime}\right),$
and by the proof of lemma 24, this expression is zero. Since $S_{t+1;f_{2},\ldots,f_{m};M}=qS_{t;f_{2},\ldots,f_{m};M}$ for $t\geq d$, it follows that the statement of the lemma holds for $t\geq d$ and $t\leq-1$.
So assume $10\leq t\leq d-$ 1. By 27, we have
$S_{t;f_{2},\ldots,f_{m};M}$ $=\frac{q^{t-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[\sum_{k=0}^{d-t-2}\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\sum_{f\in\mathcal{M}_{k}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right.$
$\qquad\left.+\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi(-\lambda)\sum_{f\in\mathcal{M}_{d-t-1}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right].$
$=\frac{q^{t-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[\sum_{\lambda\in\mathbb{F}_{q}^{\times}}\prod_{i=s+1}^{m}\left(\frac{\lambda}{f_{i}}\right)_{\chi}^{M_{1,i}}\psi(-\lambda)\sum_{f\in\mathcal{M}_{d-t-1}}a\left(f,f_{2},\ldots,f_{m};M^{\prime}\right)\right]$
$=\frac{q^{t-d}}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\left[\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}\right],$
as desired. ∎
We then obtain the following functional equation.
Proposition 33.
$P_{f_{2},\ldots,f_{m};M}(u_{1})=\frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{q\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right).$
Proof.
Indeed, we have
$u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)$ $=u_{1}^{d-1}\left(\sum_{t=0}^{d-1}S_{t;f_{2},\ldots,f_{m};M^{\prime}}q^{-t}u_{1}^{-t}+\frac{S_{d;f_{2},\ldots,f_{m};M^{\prime}}q^{-d}u_{1}^{-d}}{1-u_{1}^{-1}}\right)$
$=\sum_{t=0}^{d-1}S_{t;f_{2},\ldots,f_{m};M^{\prime}}q^{-t}u_{1}^{d-t-1}.$
So, by lemma 32 and changing variable ($t\mapsto d-t-1$), we obtain
$P_{f_{2},\ldots,f_{m};M}(u_{1})$ $=\sum_{t=0}^{d-1}\frac{q^{t-d}\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}S_{d-t-1;f_{2},\ldots,f_{m};M^{\prime}}u_{1}^{t}$
$=\frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}\sum_{t=0}^{d-1}q^{-1}q^{-t}S_{t;f_{2},\ldots,f_{m};M^{\prime}}u_{1}^{d-t-1}$
$=\frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{q\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right),$
as desired. ∎
4.3 Putting everything together
We now complete the first step of the proof of theorem 6.
Let $\zeta_{n}$ be a primitive $n$th root of unity, e.g. $\zeta_{n}=e^{2\pi i/n}$.
Recall the multiple Dirichlet series
$L\left(u_{1},\ldots,u_{m};M\right)$ $=\sum_{d_{2},\ldots,d_{m}\geq 0}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}\sum_{t\geq 0}\sum_{f\in\mathcal{M}_{t}}a\left(f,f_{2},\ldots,f_{m};M\right)u_{1}^{t}u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}$
$=\sum_{d_{2},\ldots,d_{m}\geq 0}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}$
and
$L_{\operatorname{fudge}}\left(u_{1},\ldots,u_{m};M\right)$
$=\sum_{f_{1},\ldots,f_{s}\in\mathbb{F}_{q}[t]^{+}}\sum_{d_{s+1},\ldots,d_{m}}b\left(d_{\geq s+1};M_{1,\geq s+1}\right)\sum_{f_{s+1}\in\mathcal{M}_{d_{s+1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M^{\prime}\right)u_{1}^{d_{1}}\cdots u_{m}^{d_{m}}.$
Note that the “roots-of-unity filter” picks out only the terms $u_{1}^{d_{1}}\cdots u_{m}^{d_{m}}$ such that $\sum_{i=s+1}^{m}d_{i}M_{1,i}$ is divisible by $n$:
$\frac{1}{n}\sum_{0\leq j\leq n-1}L\left(u_{1},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}u_{m};M\right)$
$=\sum_{\underset{n|\sum_{i=s+1}^{m}d_{i}M_{1,i}}{d_{2},\ldots,d_{m}}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}.$
So, by proposition 33, we have
$L\left(u_{1},\ldots,u_{m};M\right)-\frac{1}{n}\sum_{0\leq j\leq n-1}L\left(u_{1},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}u_{m};M\right)$
$=\sum_{\underset{n\nmid\sum_{i=s+1}^{m}d_{i}M_{1,i}}{d_{2},\ldots,d_{m}}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}$
$=\sum_{\underset{n\nmid\sum_{i=s+1}^{m}d_{i}M_{1,i}}{d_{2},\ldots,d_{m}}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}\frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{q\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}$
$=\frac{1}{u_{1}}\sum_{\underset{n\nmid\sum_{i=s+1}^{m}d_{i}M_{1,i}}{d_{2},\ldots,d_{m}}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}\left(\frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{q^{1+\sum_{i=s+1}^{m}d_{i}/2}\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)\right.$
$\qquad\left.\cdot\prod_{j=2}^{s}u_{j}^{d_{j}}\prod_{j=s+1}^{m}\left(q^{1/2}u_{1}u_{j}\right)^{d_{j}}\right)$
$=\frac{1}{u_{1}}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots u_{s},q^{1/2}u_{1}u_{s+1},\ldots,q^{1/2}u_{1}u_{m};M\right)$
$\qquad-\frac{1}{u_{1}}\frac{1}{n}\sum_{0\leq j\leq n-1}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}q^{1/2}u_{1}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}q^{1/2}u_{1}u_{m};M\right).$
Also, by proposition 31, we have
$\frac{1}{n}\sum_{0\leq j\leq n-1}\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}u_{m};M\right)$
$=\sum_{\underset{n|\sum_{i=s+1}^{m}d_{i}M_{1,i}}{d_{2},\ldots,d_{m}}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}(u_{1})u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}(qu_{1}-1)$
$=\sum_{\underset{n|\sum_{i=s+1}^{m}d_{i}M_{1,i}}{d_{2},\ldots,d_{m}}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}\frac{u_{1}^{d-1}\left(1-u_{1}\right)}{q^{\sum_{i=s+1}^{m}d_{i}/2}\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)\prod_{j=2}^{s}u_{j}^{d_{j}}\prod_{j=s+1}^{m}\left(q^{1/2}u_{1}u_{j}\right)^{d_{j}}$
$=\frac{1}{n}\sum_{0\leq j\leq n-1}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}q^{1/2}u_{1}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}q^{1/2}u_{1}u_{m};M\right)\frac{1-u_{1}}{u_{1}}.$
Then, combining these two computations yields
$\frac{qu_{1}-1}{u_{1}}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots u_{s},q^{1/2}u_{1}u_{s+1},\ldots,q^{1/2}u_{1}u_{m};M\right)$
$\qquad-\frac{qu_{1}-1}{u_{1}}\frac{1}{n}\sum_{0\leq j\leq n-1}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}q^{1/2}u_{1}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}q^{1/2}u_{1}u_{m};M\right)$
$=\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)-\frac{1}{n}\sum_{0\leq j\leq n-1}\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}u_{m};M\right)$
$=\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)$
$\qquad+\left(1-\frac{1}{u_{1}}\right)\frac{1}{n}\sum_{0\leq j\leq n-1}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}q^{1/2}u_{1}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}q^{1/2}u_{1}u_{m};M\right).$
Re-arranging, we finally have
$u_{1}\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)$
$=\left(qu_{1}-1\right)L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots u_{s},q^{1/2}u_{1}u_{s+1},\ldots,q^{1/2}u_{1}u_{m};M\right)$
$-\frac{qu_{1}+u_{1}-2}{n}\sum_{0\leq j\leq n-1}L_{\operatorname{fudge}}\left(\frac{1}{qu_{1}},u_{2},\ldots,u_{s},\zeta_{n}^{jM_{1,s+1}}q^{1/2}u_{1}u_{s+1},\ldots,\zeta_{n}^{jM_{1,m}}q^{1/2}u_{1}u_{m};M\right),$
which completes the proof of the functional equations as a formal equality (theorem 6).
4.4 A short verification
We explain how our functional equations match up with the functional equations in Whitehead’s thesis when our methods apply (when $M_{1,1}=0$), which, as we mentioned earlier, use the axioms from Diaconu and Pasol’s paper—this is completely analogous to how we derive functional equations using the axioms from Sawin’s paper. One key difference is that Whitehead’s argument is purely numerical, whereas our argument is a combination of numerical and geometric methods.
Let us first recast the functional equations appearing as equations (2.2.3) and (2.2.4) of Whitehead’s thesis in terms of the notation of our paper; we will freely use the notation of [whitehead_multiple] in this example.
When Whitehead writes $j\sim i$, this is equivalent to the matrix entry $M_{i,j}$ being equal to one. Otherwise, $M_{i,j}=0$. He also requires $M_{i,i}=0$ for all $i$. Hence, without loss of generality, in our setting where we require $s\geq 1$, the symmetric matrix $M$ looks like
$\begin{bmatrix}0&0&\cdots&0&1&\cdots&1\\ 0&0&*&*&*&*&*\\ \vdots&*&\ddots&*&*&*&*\\ 0&*&*&\ddots&*&*&*\\ 1&*&*&*&\ddots&*&*\\ \vdots&*&*&*&*&\ddots&*\\ 1&*&*&*&*&*&0\end{bmatrix},$
where the blank spots are arbitrary (1 or 0).
Since $\chi$ is a non-trivial quadratic character in [whitehead_multiple], we have $n=2.$
Note that $M^{\prime}=M$ because
  1. (i)
    $M^{\prime}_{i,j}=M_{i,j}+M_{1,i}+M_{1,j}=M_{i,j}+2\equiv M_{i,j}$ for $j>i\geq s+1,$
  2. (ii)
    $M^{\prime}_{i,i}=M_{i,i}+M_{1,i}+n/2=M_{i,i}+1+1\equiv M_{i,i}$ for $i\geq s+1$,
  3. (iii)
    $M^{\prime}_{1,i}=-M_{1,i}\equiv M_{1,i}$ for all $i$,
  4. (iv)
    $M_{i,j}^{\prime}=M_{i,j}$ for $j>i>1$ and $i\leq s$, and
  5. (v)
    $M_{i,i}^{\prime}=M_{i,i}$ for $i\leq s$.
In Whitehead’s notation, $a_{i}=d_{i},$ $x_{i}=u_{i},$ and there are two functional equations depending on the parity of $\sum_{j\sim 1}a_{j}$ (note that Whitehead establishes functional equations for every $i$, but we can only do so for $i\in\left\{1,\ldots,s\right\}$).
Since $q\equiv 1\bmod 4$ (by assumption in [whitehead_multiple]), we have
$\operatorname{fudge}\left(d_{2},\ldots,d_{m};M\right)$ $=\frac{\chi(-1)^{\sum_{s+1\leq i<j\leq m}d_{i}d_{j}M_{1,i}}(-1)^{\sum_{i\geq s+1}\frac{d_{i}(d_{i}-1)(q-1)}{4}}}{\prod_{i\geq s+1}G\left(\chi^{M_{1,i}},\psi\right)^{d_{i}}}$
$=\frac{1}{G\left(\chi,\psi\right)^{\sum_{i\geq s+1}d_{i}}}$
$=\frac{1}{\left(q^{1/2}\right)^{\sum_{i\geq s+1}d_{i}}}$
where we use the fact that the unique non-trivial quadratic character $\chi\colon\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ is given by $\left(\frac{\operatorname{Nm}_{\mathbb{F}_{q}/\mathbb{F}_{p}}\left(-\right)}{p}\right)$, which means $\chi(-1)=1$ (if $p\equiv 1\bmod 4$, then $\left(\frac{-1}{p}\right)=1$ by quadratic reciprocity; if $p\equiv 3\bmod 4,$ then $q$ is necessarily an even power of $p$), as well as Gauss’s computation of a quadratic Gauss sum.
The case $\sum_{j\sim 1}a_{j}$ is odd means $\sum_{i\geq s+1}d_{i}$ is odd, i.e. the case where $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}}$ is not trivial on $\mathbb{F}_{q}^{\times}$ from subsection 4.1.
Then, proposition 33 tells us that
$P_{f_{2},\ldots,f_{m};M}(u_{1})$ $=\frac{\chi(-1)^{\sum_{i=s+1}^{m}d_{i}M_{1,i}}G\left(\chi^{\sum_{i=s+1}^{m}d_{i}M_{1,i}},\psi\right)}{q\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)$
$=\frac{G\left(\chi,\psi\right)\left(q^{1/2}\right)^{\sum_{i\geq s+1}d_{i}}}{q}u_{1}^{d-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)$
$=\left(q^{1/2}u_{1}\right){}^{\left(\sum_{i\geq s}d_{i}\right)-1}P_{f_{2},\ldots,f_{m};M^{\prime}}\left(q^{-1}u_{1}^{-1}\right).$
Summing both sides over all $f_{i}\in\mathcal{M}_{d_{i}}$ for $i\in\left\{2,\ldots,m\right\}$ gives the functional equation (2.2.3) of [whitehead_multiple].
The case $\sum_{j\sim i}a_{j}$ is even means $\sum_{i\geq s}d_{i}$ is even, i.e. the first case where $\prod_{i=s+1}^{m}\left(\frac{-}{f_{i}}\right)_{\chi}^{M_{1,i}}$ is trivial on $\mathbb{F}_{q}^{\times}$. Then, we know that
$P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)\left(qu_{1}-1\right)$ $=\frac{1}{\operatorname{fudge}\left(d_{\geq s+1};M_{1,\geq s+1}\right)}u_{1}^{d-1}\left(1-u_{1}\right)P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right)$
$=\left(q^{1/2}u_{1}\right){}^{\sum_{i\geq s}d_{i}}\left(\frac{1-u_{1}}{u_{1}}\right)P_{f_{2},\ldots,f_{m};M^{\prime}}\left(\frac{1}{qu_{1}}\right),$
so
$\left(1-qu_{1}\right)P_{f_{2},\ldots,f_{m};M}(u_{1})=\left(q^{1/2}u_{1}\right){}^{\sum_{i\geq s+1}d_{i}}\left(1-u_{1}^{-1}\right)P_{f_{2},\ldots,f_{m};M^{\prime}}\left(q^{-1}u_{1}^{-1}\right).$
Summing both sides over all $f_{i}\in\mathcal{M}_{d_{i}}$ for $i\in\left\{2,\ldots,m\right\}$ gives the functional equation (2.2.5) of [whitehead_multiple].
5   Bounds on $a$-coefficients and their sums
We now begin the second step of the proofs of the main results. Using the notation in [sawin_general], along with notation of the previous section, let
$\lambda\left(d_{1},\ldots,d_{m};M\right)=\sum_{f_{1}\in\mathcal{M}_{d_{1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M\right).$
In this section, we obtain upper bounds for $a\left(f_{1},\ldots,f_{m};M\right)$ and $\lambda\left(d_{1},\ldots,d_{m};M\right)$. To do so, we use the Grothendieck-Lefschetz fixed point formula to bound $\lambda\left(d_{1},\ldots,d_{m};M\right)$ in terms of traces of Frobenius acting on compactly-supported cohomology groups with coefficients in $K_{d_{1},\ldots,d_{m}}$. These, in turn, can be bounded in terms of the dimensions of the cohomology groups. These cohomology groups can be viewed as direct summands of cohomology groups of a suitable compactification (via Kontsevich moduli spaces of stable maps) with coefficients in a lisse rank one sheaf using the decomposition theorem for perverse sheaves. Finally, we can use the earlier results of subsection 2.4.
To bound $a\left(f_{1},\ldots,f_{m};M\right)$, we use the bound for $\lambda\left(d_{1},\ldots,d_{m};M\right)$ along with the axioms of a multiple Dirichlet series, namely the local-to-global relationship and normalization (the last four axioms).
Let $r=d_{1}+\cdots+d_{m}$.
By the Grothendieck-Lefschetz fixed point formula (c.f. [Fu]), we have
$\lambda\left(d_{1},\ldots,d_{m};M\right)$ $=\sum_{f_{1}\in\mathcal{M}_{d_{1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M\right)$
$=\sum_{f_{1}\in\mathcal{M}_{d_{1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}\sum_{i}\left(-1\right)^{i}\operatorname{Tr}\left(\operatorname{Fr}_{q},\mathcal{H}^{i}\left(K_{d_{1},\ldots,d_{m}}\right)_{\left(f_{1},\ldots,f_{m}\right)}\right)$
$=\sum_{i}\left(-1\right)^{i}\operatorname{Tr}\left(\operatorname{Fr}_{q},H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)\right).$
Then, using the fact that $K_{d_{1},\ldots,d_{m}}$ is pure of weight zero and Artin vanishing (c.f. [Fu]), we obtain
$\left|\lambda\left(d_{1},\ldots,d_{m};M\right)\right|$ $\leq\sum_{i=0}^{2r}\left|\operatorname{Tr}\left(\operatorname{Fr}_{q},H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)\right)\right|$
$\leq\sum_{i=0}^{2r}\left(\dim H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)\right)q^{i/2}.$ (5.1)
Note that the quotient map $\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}\to\left[\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]$ is proper and etale, and consequently $p\colon\left[\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\right]\to\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\cong\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}$ is proper and quasi-finite. Let $X$ denote the locus of tuples $\left(f_{1},\ldots,f_{m}\right)$ such that $\left(f_{i},f_{j}\right)=\left(f_{i},f_{i}^{\prime}\right)=1$ for $i\neq j$ for all $1\leq i,j\leq m$, i.e. the pure configuration space of $r$ points on $\mathbb{A}^{1}$ modulo $S_{d_{1}}\times\cdots\times S_{d_{m}}.$ Consider the following commutative diagram:
[…]
By the decomposition theorem for perverse sheaves from [bbdg], we have
$K_{d_{1},\ldots,d_{m}}$ $=j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\left[-r\right]$
$\inplus p_{*}\left(j_{\text{stack}}\right)_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\left[-r\right],$
which means
$H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)$ $\inplus H_{c}^{i}\left(\left[\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],\left(j_{\text{stack}}\right)_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\left[-r\right]\right)$
$=H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\left[-r\right]\right)^{S_{d_{1}}\times\cdots\times S_{d_{m}}},$ (5.2)
where by abuse of notation $j\colon X^{\prime}\subset\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}$ is the open immersion of the pure configuration space of $r$ distinct points on $\mathbb{A}^{1}$ into $\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}}$.
[…]
Lemma 34.
$\operatorname{ev}\colon\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{X^{\prime}}\to X^{\prime}$ is an isomorphism.
Proof.
¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ is given by chains of $\mathbb{P}^{1}$’s, along with a distinguished $\mathbb{P}^{1}$ that maps isomorphically (of degree 1) onto $\mathbb{P}^{1}$. Moreover, every $\mathbb{P}^{1}$ that is not the single distinguished $\mathbb{P}^{1}$ must be stable, i.e. have at least three special points. But by assumption on $X$, this is only possible if the twig is a single copy of $\mathbb{P}^{1}$, from which the result follows. ∎
Let the open immersion ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{X^{\prime}}\cong X^{\prime}\subset\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ be denoted by $g$.
Note that $\operatorname{ev}$ is proper because ¯$\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)$ and $\left(\mathbb{P}^{1}\right)^{r}$ are proper.
Lemma 35.
$j^{*}R\operatorname{ev}_{*}g_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\cong\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]$.
Proof.
By proper base change (c.f. [Fu]), it is evident that the left-hand side is isomorphic to $R\operatorname{ev}_{*}g^{*}g_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)$, so the result follows from lemma 34. ∎
By the decomposition theorem applied to the proper map $\operatorname{ev}$ and lemma 35, it follows that $R\operatorname{ev}_{*}g_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)$ contains $j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)$ as a direct summand.
As a result, we have
$H_{c}^{i}\left(\mathbb{A}^{r},j_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\right)\inplus H_{c}^{i}\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}},g_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\right).$ (5.3)
Hence, combining (5.2) and (5.3) after taking $S_{d_{1}}\times\cdots\times S_{d_{m}}$-invariants, we obtain
$\dim H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)\leq\dim H_{c}^{i}\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}},g_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)\left[-r\right]\right)^{S_{d_{1}}\times\cdots\times S_{d_{m}}}.$ (5.4)
[…]
$g_{!*}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\left[r\right]\right)=g_{!}^{\prime}g_{*}^{\prime\prime}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\right)\left[r\right],$
[…]
Hence, (5.4) can be written as
$\dim H_{c}^{i}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)\leq\dim H_{c}^{i}\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}},g_{!}^{\prime}g_{*}^{\prime\prime}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\right)\right)^{S_{d_{1}}\times\cdots\times S_{d_{m}}}.$ (5.5)
[EVW] prove  [^1] Their statement is for $L$ a constant sheaf, but the proof is identical. the following (as proposition 7.7):
Proposition 36.
Let $A$ be a Henselian discrete valuation ring whose quotient field has characteristic zero. Let $\overline{\eta}$ and $\overline{s}$ be the generic and special points of $\operatorname{Spec}A$, respectively.
Suppose $X$ is a scheme smooth and proper over $\operatorname{Spec}A$, $D\hookrightarrow X$ is a normal crossings divisor relative to $\operatorname{Spec}A$, and $U=X\backslash D$. Let $G$ be a finite group acting on $X$ and $U$ compatibly and $L$ a lisse sheaf on $U$.
Then,
$H_{c}^{i}\left(U_{\overline{\eta}},L\right)\cong H_{c}^{i}\left(U_{\overline{s}},L\right)$
as $G$-modules for all $i$.
Let us apply this proposition to our situation. By setting $A$ to be the Witt vectors $W\left(\mathbb{F}_{p}\right)$, and since $g_{*}^{\prime\prime}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\right)$ is a lisse sheaf (by proposition 10), proposition 36 tells us that we may actually work in the characteristic zero situation.
To bound
$\dim H_{c}^{i}\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}},g_{!}^{\prime}g_{*}^{\prime\prime}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\right)\right)^{S_{d_{1}}\times\cdots\times S_{d_{m}}}$
$=\dim H_{c}^{i}\left(\left[\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],g_{!}^{\prime}g_{*}^{\prime\prime}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\right)\right)$
$=\dim H_{c}^{i}\left(\left[\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}\backslash D\right)/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],g_{*}^{\prime\prime}\left(\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)\right)\right),$
where $D$ is the union of divisors for which the local monodromy of $\mathcal{L}_{\chi}\left(F_{d_{1},\ldots,d_{m}}\right)$ is non-trivial, Poincare duality (c.f. [Fu]) implies that it suffices to bound the more general quantity
$\dim H^{*}\left(\left[\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}\backslash D\right)/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],\mathcal{L}\right),$
for any rank one lisse sheaf $\mathcal{L}$ on [(¯$\left[\left(\overline{\mathcal{M}}_{0,r}\left(\mathbb{P}^{1},1\right)_{\mathbb{A}^{r}}\backslash D\right)/S_{d_{1}}\times\cdots\times S_{d_{m}}\right],$ which we already did in subsection 2.4.
Let $C=64m$. Combining (5.5), proposition 36, and proposition 17, we obtain the bound
$\dim H_{c}^{*}\left(\prod_{j=1}^{m}\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)\ll_{m}\left(Cm\right)^{r}.$ (5.6)
Applying this to (5.1), we get the following:
Proposition 37.
$\left|\lambda\left(d_{1},\ldots,d_{m};M\right)\right|$ $\ll_{m}\left(Cm\right)^{r}\sum_{i=0}^{2r}q^{i/2}$
$\ll_{q,m}\left(Cmq\right)^{r}.$
Let us now bound $a\left(f_{1},\ldots,f_{m};M\right)$ for $f_{i}\in\mathcal{M}_{d_{i}}$.
Note that the fourth axiom of theorem 9 implies that
$\lambda\left(d_{1},\ldots,d_{m};M\right)$ $=\sum_{f_{1}\in\mathcal{M}_{d_{1}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}a\left(f_{1},\ldots,f_{m};M\right)$
$=\sum_{j\in J\left(d_{1},\ldots,d_{m};q,\chi,M\right)}c_{j}\cdot\frac{q^{r}}{\overline{\alpha_{j}}}.$
By theorem 9, $c_{j}$ is the signed multiplicity of $\alpha_{j},$ which is an eigenvalue of $\operatorname{Fr}_{q}$ acting on the complex $\left(K_{d_{1},\ldots,d_{m}}\right)_{\left(t^{d_{1}},\ldots,t^{d_{m}}\right)}$. By $\mathbb{G}_{m}$-localization (lemma 2.16 of [sawin_general]), we have
$H_{c}^{i}\left(\prod\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right)=\mathcal{H}^{i}\left(\left(K_{d_{1},\ldots,d_{m}}\right)_{\left(T^{d_{1}},\ldots,T^{d_{m}}\right)}\right),$
from which it follows that $\max_{j\in J\left(d_{1},\ldots,d_{m};q,\chi,M\right)}c_{j}\leq\max_{i}\dim H_{c}^{i}\left(\prod\mathbb{A}_{\overline{\mathbb{F}_{q}}}^{d_{j}},K_{d_{1},\ldots,d_{m}}\right).$ Then, (5.6) gives
$c_{j}\ll_{m}C^{d_{1}+\cdots+d_{m}}.$
This implies the following bound on $a$-coefficients:
Corollary 38.
Let $\pi$ be a prime. Then, for $\sum_{i}e_{i}=0$ or 1, we have $\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}}\right)\right|=1$, and for $\sum_{i}e_{i}\geq 2$, we have
$\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}}\right)\right|\ll_{m}C^{\sum_{i}e_{i}}q^{-\deg\pi}q^{\frac{\sum_{i}e_{i}}{2}\deg\pi}.$
Proof.
If $\sum_{i}e_{i}=0$ or 1, $K_{e_{1},\ldots,e_{m}}$ is the constant sheaf. Then, by the third axiom of theorem 9, we have
$\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}}\right)\right|=\left|\sum_{j\in J\left(e_{1},\ldots,e_{m};q,\chi,M\right)}c_{j}\alpha_{j}^{\deg\pi}\right|=1.$
For $\sum_{i}e_{i}\geq 2$, we have
$\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}}\right)\right|$ $=\left|\sum_{j\in J\left(e_{1},\ldots,e_{m};q,\chi,M\right)}c_{j}\alpha_{j}^{\deg\pi}\right|$
$\leq\sup_{j\in J\left(e_{1},\ldots,e_{m};q,\chi,M\right)}\left|c_{j}\right|\left(q^{\frac{\sum_{i}e_{i}}{2}-1}\right)^{\deg\pi}$
$\leq C^{\sum_{i}e_{i}}\left(q^{\frac{\sum_{i}e_{i}}{2}-1}\right)^{\deg\pi}$
$=C^{\sum_{i}e_{i}}q^{-\deg\pi}q^{\frac{\sum_{i}e_{i}}{2}\deg\pi},$
where in the second step we use the fifth axiom of theorem 9. ∎
6   Meromorphic continuation of multiple Dirichlet series
Using the bounds of the previous section, we are finally able to finish the proofs of theorems 1 and 6.
Again, let us continue using the notation of the previous section. In particular, recall that $C=64m$.
First, we show the following:
Proposition 39.
$\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)$ and $\left(qu_{1}-1\right)L_{\operatorname{fudge}}\left(u_{1},\ldots,u_{m};M\right)$ converge for
$\left|u_{s+1}\right|,\ldots,\left|u_{m}\right|<\frac{1}{Cq\max\left\{Cq\left|u_{1}\right|,1\right\}};\left|u_{2}\right|,\ldots,\left|u_{s}\right|<\frac{1}{Cq}.$
Remark 40.
Note that the inequalities
$\left|u_{s+1}\right|,\ldots,\left|u_{m}\right|<\frac{1}{Cq\max\left\{Cq\left|u_{1}\right|,1\right\}};\left|u_{2}\right|,\ldots,\left|u_{s}\right|<\frac{1}{Cq}$
and
$\left|q^{1/2}u_{1}u_{s+1}\right|,\ldots,\left|q^{1/2}u_{1}u_{m}\right|<\frac{1}{Cq\max\left\{Cq\left|\frac{1}{qu_{1}}\right|,1\right\}};\left|u_{2}\right|,\ldots,\left|u_{s}\right|<\frac{1}{Cq}$
are simultaneously satisfied by
$u_{1}\neq 0;\left|u_{1}\right|,\ldots,\left|u_{s}\right|<\frac{1}{Cq};\left|u_{s+1}\right|,\ldots,\left|u_{m}\right|<\frac{1}{C^{2}q^{3/2}};\left|u_{2}\right|,\ldots,\left|u_{s}\right|<\frac{1}{Cq}.$
Hence, the proposition completes the proofs of theorems 1 and 6.
Proof of proposition.
Let us prove convergence for $\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)$ first.
From lemma 28, we have
$\left(qu_{1}-1\right)P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)=-S_{d_{s+1}+\cdots+d_{m};f_{2},\ldots,f_{m};M}u_{1}^{d_{s+1}+\cdots+d_{m}}+\sum_{t=0}^{d_{s+1}+\cdots+d_{m}-1}\left(qu_{1}-1\right)S_{t;f_{2},\ldots,f_{m};M}u_{1}^{t}.$
Summing over all $f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}$, we obtain
$\left(qu_{1}-1\right)\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)$
$=\sum_{t=0}^{d_{s+1}+\cdots+d_{m}}\left(q\lambda\left(t-1,d_{2},\ldots,d_{m};M\right)-\lambda\left(t,d_{2},\ldots,d_{m};M\right)\right)u_{1}^{t}.$
By proposition 37, we have
$\left|q\lambda\left(t-1,d_{2},\ldots,d_{m};M\right)-\lambda\left(t,d_{2},\ldots,d_{m};M\right)\right|$
$\leq q\left|\lambda\left(t-1,d_{2},\ldots,d_{m};M\right)\right|+\left|\lambda\left(t,d_{2},\ldots,d_{m};M\right)\right|$
$\ll_{q,m}\left(Cq\right)^{t+d_{2}+\cdots+d_{m}}.$
It follows that
$\left|\left(qu_{1}-1\right)\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}\left(u_{1}\right)\right|$
$\leq\sum_{t=0}^{d_{s+1}+\cdots+d_{m}}\left|q\lambda\left(t-1,d_{2},\ldots,d_{m};M\right)-\lambda\left(t,d_{2},\ldots,d_{m};M\right)\right|\left|u_{1}\right|^{t}$
$\ll_{q,m}\sum_{t=0}^{d_{s+1}+\cdots+d_{m}}\left(Cq\right)^{d_{2}+\cdots+d_{m}}\left|Cqu_{1}\right|^{t}$
$\ll_{q,m}\left(Cq\right)^{d_{2}+\cdots+d_{m}}\left(1+d_{s+1}+\cdots+d_{m}\right)\max\left\{1,\left|Cqu_{1}\right|^{d_{s+1}+\cdots+d_{m}}\right\}.$
Now, recall
$L\left(u_{1},\ldots,u_{m};M\right)=\sum_{d_{2},\ldots,d_{m}}\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}(u_{1})u_{2}^{d_{2}}\cdots u_{m}^{d_{m}},$
so
$\left|\left(qu_{1}-1\right)L\left(u_{1},\ldots,u_{m};M\right)\right|$
$\leq\sum_{d_{2},\ldots,d_{m}}\left|\left(qu_{1}-1\right)\sum_{f_{2}\in\mathcal{M}_{d_{2}},\ldots,f_{m}\in\mathcal{M}_{d_{m}}}P_{f_{2},\ldots,f_{m};M}(u_{1})u_{2}^{d_{2}}\cdots u_{m}^{d_{m}}\right|$
$\ll_{q,m}\sum_{d_{2},\ldots,d_{m}}\left(Cq\right)^{d_{2}+\cdots+d_{m}}\left(1+d_{s+1}+\cdots+d_{m}\right)\max\left\{1,\left|Cqu_{1}\right|^{d_{s+1}+\cdots+d_{m}}\right\}\left|u_{2}\right|^{d_{2}}\cdots\left|u_{m}\right|^{d_{m}}$
$=\sum_{d_{2},\ldots,d_{m}}\left(1+d_{s+1}+\cdots+d_{m}\right)\max\left\{1,\left|Cqu_{1}\right|^{d_{s+1}+\cdots+d_{m}}\right\}\left|Cqu_{2}\right|^{d_{2}}\cdots\left|Cqu_{m}\right|^{d_{m}}$
$=\sum_{d_{2},\ldots,d_{m}}\left(1+d_{s+1}+\cdots+d_{m}\right)\left|Cqu_{2}\right|^{d_{2}}\cdots\left|Cqu_{s}\right|^{d_{s}}\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{s+1}\right|^{d_{s+1}}$
$\qquad\cdots\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{m}\right|^{d_{m}}$
$=\left(\prod_{i=2}^{s}\frac{1}{1-Cq\left|u_{i}\right|}\right)\sum_{d_{s+1},\ldots,d_{m}}\left(1+d_{s+1}+\cdots+d_{m}\right)\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{s+1}\right|^{d_{s+1}}$
$\qquad\cdots\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{m}\right|^{d_{m}},$
where in the last step we use the assumption that $\left|u_{2}\right|,\ldots,\left|u_{s}\right|<\frac{1}{Cq}$.
It suffices to show
$\sum_{d_{s+1},\ldots,d_{m}}\left(1+d_{s+1}+\cdots+d_{m}\right)\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{s+1}\right|^{d_{s+1}}\cdots\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{m}\right|^{d_{m}}$
converges. To see this, we can write the expression as
$\sum_{d_{s+1},\ldots,d_{m}}\prod_{j=s+1}^{m}\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{j}\right|^{d_{j}}+\sum_{i=s+1}^{m}\sum_{d_{s+1},\ldots,d_{m}}d_{i}\prod_{j=s+1}^{m}\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{j}\right|^{d_{j}}$
$=\prod_{j=s+1}^{m}\frac{1}{1-\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{j}\right|}$
$\qquad+\sum_{i=s+1}^{m}\frac{\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{i}\right|}{1-\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{i}\right|}\prod_{j=s+1}^{m}\frac{1}{1-\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{j}\right|},$
which evidently converges for $\left|\max\left\{1,\left|Cqu_{1}\right|\right\}Cqu_{j}\right|<1$ for $j\in\left\{s+1,\ldots,m\right\}$.
For $L_{\operatorname{fudge}}(u_{1},\ldots,u_{m};M)$, note that the argument for $L\left(u_{1},\ldots,u_{m};M\right)$ works without any modification since $\left|b\left(d_{\geq s+1};M_{1,\geq s+1}\right)\right|$ is bounded independently of $d_{\geq s+1}$. ∎
Finally, we demonstrate how to obtain a different region of convergence using the bounds on $a$-coefficients. In particular, this region neither contains nor is contained in the region defined in proposition 39, so meromorphic continuation allows us to extend the domain of definition of the multiple Dirichlet series.
Proposition 41.
$L\left(u_{1},\ldots,u_{m};M\right)$ and $L_{\operatorname{fudge}}(u_{1},\ldots,u_{m};M)$ converge for
$\left|u_{i}\right|<\min\left\{q^{-1},C^{-1}q^{-1/2}\right\}.$
Proof.
Let us prove convergence for $L\left(u_{1},\ldots,u_{m};M\right)$ first.
To get bounds on radii of convergence, it suffices to obtain radii of convergence for the Euler product
$\prod_{\pi\text{ prime}}\left(\sum_{e_{1},\ldots,e_{m}}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|u_{1}^{e_{1}\deg\pi}\cdots u_{m}^{e_{m}\deg\pi}\right).$
Indeed, observe that showing convergence for
$L\left(u_{1},\ldots,u_{m};M\right)=\sum_{f_{1},\ldots,f_{m}\in\mathbb{F}_{q}[t]^{+}}a\left(f_{1},\ldots,f_{m};M\right)u_{1}^{\deg f_{1}}\cdots u_{m}^{\deg f_{m}}$
is implied by convergence for
$\sum_{f_{1},\ldots,f_{m}\in\mathbb{F}_{q}[t]^{+}}\left|a\left(f_{1},\ldots,f_{m};M\right)\right|u_{1}^{\deg f_{1}}\cdots u_{m}^{\deg f_{m}},$
and twisted multiplicativity gives
$\left|a\left(f_{1},\ldots,f_{m};M\right)\right|=\prod_{\pi\text{ prime}}\left|a\left(\pi^{v_{\pi}\left(f_{1}\right)},\ldots,\pi^{v_{\pi}\left(f_{m}\right)}\right)\right|.$
By corollary 38, the inner term
$\sum_{e_{1},\ldots,e_{m}}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}$
$=\sum_{\left(e_{1},e_{2},\ldots,e_{m}\right)=0}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}$
$\qquad+\sum_{e_{1}+\cdots+e_{m}=1}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}$
$\qquad+\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)=0}{e_{1}\geq 2}}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}$
$\qquad+\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}$
$\leq 1+\left|u_{1}\right|^{\deg\pi}+\cdots+\left|u_{m}\right|^{\deg\pi}+\sum_{e_{1}\geq 2}\left|u_{1}\right|^{e_{1}\deg\pi}$
$\qquad+\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}C^{\sum_{i}e_{i}}q^{-\deg\pi}q^{\frac{\sum_{i}e_{i}}{2}\deg\pi}\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}.$
To establish absolute convergence of
$\prod_{\pi\text{ prime}}\left(\sum_{e_{1},\ldots,e_{m}}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|u_{1}^{e_{1}\deg\pi}\cdots u_{m}^{e_{m}\deg\pi}\right),$
it suffices (and is equivalent) to establishing absolute convergence of
$\sum_{\pi\text{ prime}}\left(\sum_{e_{1},\ldots,e_{m}}\left|a\left(\pi^{e_{1}},\ldots,\pi^{e_{m}};M\right)\right|u_{1}^{e_{1}\deg\pi}\cdots u_{m}^{e_{m}\deg\pi}-1\right).$
By the above, this is at most
$\leq\sum_{\pi\text{ prime}}\left(\left|u_{1}\right|^{\deg\pi}+\cdots+\left|u_{m}\right|^{\deg\pi}+\sum_{e_{1}\geq 2}\left|u_{1}\right|^{e_{1}\deg\pi}\right.$
$\qquad\left.+q^{-\deg\pi}\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}C^{\sum_{i}e_{i}}q^{\frac{\sum_{i}e_{i}}{2}\deg\pi}\left|u_{1}\right|^{e_{1}\deg\pi}\cdots\left|u_{m}\right|^{e_{m}\deg\pi}\right)$
$\leq\sum_{d\geq 1}\left(q^{d}\left|u_{1}\right|^{d}+\cdots+q^{d}\left|u_{m}\right|^{d}+\sum_{e_{1}\geq 2}\left(q\left|u_{1}\right|^{e_{1}}\right)^{d}\right.$
$\qquad\left.+\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}C^{\sum_{i}e_{i}}q^{\frac{\sum_{i}e_{i}}{2}d}\left|u_{1}\right|^{e_{1}d}\cdots\left|u_{m}\right|^{e_{m}d}\right)$
$=\sum_{d\geq 1}\left(q|u_{1}|\right)^{d}+\cdots+\sum_{d\geq 1}\left(q|u_{m}|\right)^{d}+\sum_{d\geq 1}\sum_{e_{1}\geq 2}\left(q\left|u_{1}\right|^{e_{1}}\right)^{d}$
$\qquad+\sum_{d\geq 1}\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}C^{\sum_{i}e_{i}}q^{\frac{\sum_{i}e_{i}}{2}d}\left|u_{1}\right|^{e_{1}d}\cdots\left|u_{m}\right|^{e_{m}d}.$
First, note that
$\sum_{d\geq 1}\left(q|u_{i}|\right)^{d}=\frac{q|u_{i}|}{1-q|u_{i}|}$
converges because $\left|u_{i}\right|<1/q$.
Next,
$\sum_{d\geq 1}\sum_{e_{1}\geq 2}\left(q\left|u_{1}\right|^{e_{1}}\right)^{d}=\sum_{e_{1}\geq 2}\frac{q\left|u_{1}\right|^{e_{1}}}{1-q\left|u_{1}\right|^{e_{1}}}$
because $\left|u_{i}\right|<1/q$, and this series moreover converges absolutely because of the ratio test (c.f. [d'Alembert]):
$\frac{\frac{q\left|u_{1}\right|^{e_{1}+1}}{1-q\left|u_{1}\right|^{e_{1}+1}}}{\frac{q\left|u_{1}\right|^{e_{1}}}{1-q\left|u_{1}\right|^{e_{1}}}}=\frac{\left|u_{1}\right|\left(1-q\left|u_{1}\right|^{e_{1}}\right)}{\left(1-q\left|u_{1}\right|^{e_{1}+1}\right)}\to\left|u_{1}\right|<1.$
Finally,
$\sum_{d\geq 1}\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}C^{\sum_{i}e_{i}}q^{\frac{\sum_{i}e_{i}}{2}d}\left|u_{1}\right|^{e_{1}d}\cdots\left|u_{m}\right|^{e_{m}d}$ $=\sum_{\underset{\left(e_{2},\ldots,e_{m}\right)\neq 0}{e_{1}+\cdots+e_{m}\geq 2}}C^{\sum_{i}e_{i}}\frac{q^{\frac{\sum_{i}e_{i}}{2}}\left|u_{1}\right|^{e_{1}}\cdots\left|u_{m}\right|^{e_{m}}}{1-q^{\frac{\sum_{i}e_{i}}{2}}\left|u_{1}\right|^{e_{1}}\cdots\left|u_{m}\right|^{e_{m}}}$
converges because $\left|u_{i}\right|<1/q$. This is moreover at most
$\sum_{e\geq 2}{e+m-1\choose m-1}C^{e}\frac{q^{e/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e}}{1-q^{e/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e}},$
which converges absolutely because of the ratio test:
($\frac{{e+m\choose m-1}C^{e+1}\frac{q^{(e+1)/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e+1}}{1-q^{(e+1)/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e+1}}}{{e+m-1\choose m-1}C^{e}\frac{q^{e/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e}}{1-q^{e/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e}}}$ $=\frac{e+m}{e}Cq^{1/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}\frac{1-q^{e/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e}}{1-q^{(e+1)/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}^{e+1}}$
$\to Cq^{1/2}\max_{1\leq i\leq m}\left\{\left|u_{i}\right|\right\}$
$<1,$
because $\left|u_{i}\right|<1/q$ for the second step and $\left|u_{i}\right|<C^{-1}q^{-1/2}$ for the last step.
For $L_{\operatorname{fudge}}(u_{1},\ldots,u_{m};M)$, note that the argument for $L\left(u_{1},\ldots,u_{m};M\right)$ works without any modification since $\left|b\left(d_{\geq s+1};M_{1,\geq s+1}\right)\right|$ is bounded independently of $d_{\geq s+1}$. ∎
[…]
[heading=bibintoc]